### v1

# Quadratic projectable Runge–Kutta methods

Abstract
Runge–Kutta methods are affine equivariant: applying a method before or after an affine change of variables yields the same numerical trajectory. However, for some applications, one would like to perform numerical integration after a quadratic change of variables. For example, in Lie–Poisson reduction, a quadratic transformation reduces the number of variables in a Hamiltonian system, yielding a more efficient representation of the dynamics. Unfortunately, directly applying a symplectic Runge–Kutta method to the reduced system generally does not preserve its Hamiltonian structure, so many proposed techniques require computing numerical trajectories of the original, unreduced system.
In this paper, we study when a Runge–Kutta method in the original variables descends to a numerical integrator expressible entirely in terms of the quadratically transformed variables. In particular, we show that symplectic diagonally implicit Runge–Kutta (SyDIRK) methods, applied to a quadratic projectable vector field, are precisely the Runge–Kutta methods that descend to a method (generally not of Runge–Kutta type) in the projected variables. We illustrate our results with several examples in both conservative and non-conservative dynamics.
1 Introduction
Given an autonomous ordinary differential equation $\dot{y}=f(y)$, the chain rule implies that a $C^{1}$ observable $z=F(y)$ evolves according to $\dot{z}=F^{\prime}(y)f(y)$. An $s$-stage Runge–Kutta method for $\dot{y}=f(y)$, with time step size $h>0$, has the form
$Y_{i}$ $=y_{0}+h\sum_{j=1}^{s}a_{ij}f(Y_{j}),$ (1a)
$y_{1}$ $=y_{0}+h\sum_{i=1}^{s}b_{i}f(Y_{i}),$ (1b)
where $a_{ij}$ and $b_{i}$ are given coefficients defining the method and $Y_{i}$ are the internal stages. McLachlan and Stern [1] recently showed that, if the Runge–Kutta method preserves quadratic invariants, and if $F$ is quadratic but not necessarily an invariant of $f$, then the numerical evolution of this observable is given by
$F(y_{1})=F(y_{0})+h\sum_{i=1}^{s}b_{i}F^{\prime}(Y_{i})f(Y_{i}).$ (2)
This property is called quadratic functional equivariance. In other words, $z_{1}=F(y_{1})$ may be obtained by applying the Runge–Kutta method to the augmented system
$\dot{y}=f(y),\qquad\dot{z}=F^{\prime}(y)f(y),$
with initial condition $z_{0}=F(y_{0})$.
Suppose now that the vector field $f$ is $F$-projectable, meaning that there exists a vector field $g$ such that $F^{\prime}(y)f(y)=g\bigl{(}F(y)\bigr{)}$. We pose the question: when is it possible to define a numerical integrator for $\dot{z}=g(z)$ that agrees with the Runge–Kutta trajectory (2)? We would like such a method to be defined in terms of $z$ alone, without needing to compute a trajectory and internal stages for $y$. In this case, we say that the Runge–Kutta method descends to an integrator for $\dot{z}=g(z)$. When this holds for all quadratic $F$ and all $F$-projectable vector fields $f$, we say that the Runge–Kutta method is quadratic projectable. Restated in these terms, the question above asks: which Runge–Kutta methods are quadratic projectable?
Some recent results may be understood as answering special cases of this question. da Silva and Lessig [2] proved that the isospectral symplectic Runge–Kutta (IsoSyRK) schemes of Modin and Viviani [3] descend to Lie–Poisson integrators on quadratic Lie algebras, with intermediate stages also in the Lie algebra, if the underlying Runge–Kutta scheme is a symplectic diagonally implicit Runge–Kutta (SyDIRK) method. This is in line with previous results of Viviani [4], in which the implicit midpoint method, a particular case of SyDIRK, is shown to descend.
In this paper, we show that the class of quadratic projectable Runge–Kutta methods is precisely the class of SyDIRK methods. Our primary focus is on Lie–Poisson reduction, where $F$ is a quadratic momentum map and SyDIRK methods descend to Lie–Poisson methods in the reduced variables. However, we also analyze more general instances of quadratic projectable vector fields, including non-conservative examples such as the semidiscretized 2D Navier–Stokes equations, for which SyDIRK methods also descend to methods in the transformed variables alone.
2 Quadratic projectable methods and their descendents
In this section, we characterize quadratic projectable Runge–Kutta methods as SyDIRK methods, and we describe the methods to which they descend in the projected variables.
Let $Y$ and $Z$ be finite-dimensional [^1] We make this assumption for simplicity, and because the applications of interest are generally finite-dimensional. However, many of the results may be adapted to infinite-dimensional Banach spaces. real or complex vector spaces and $F\colon Y\rightarrow Z$ be a quadratic map, meaning that we can expand $F(y)$ around any $y_{0}\in Y$ as
$F(y)=F(y_{0})+F^{\prime}(y_{0})(y-y_{0})+\frac{1}{2}F^{\prime\prime}(y-y_{0},y-y_{0}).$
The following useful lemma describes how $F(y)\in Z$ evolves numerically whenever a Runge–Kutta method is applied to $y\in Y$.
Lemma 1.
Suppose the Runge–Kutta method (1) is applied to integrate $\dot{y}=f(y)$ on $Y$. If $F\colon Y\rightarrow Z$ is quadratic, then
$F(Y_{i})$ $=F(y_{0})$ $+h\sum_{j=1}^{s}a_{ij}F^{\prime}(Y_{j})f(Y_{j})$
$+\frac{h^{2}}{2}\sum_{j,k=1}^{s}(a_{ij}a_{ik}-a_{ij}a_{jk}-a_{ik}a_{kj})F^{\prime\prime}\bigl{(}f(Y_{j}),f(Y_{k})\bigr{)},$ (3a)
$F(y_{1})$ $=F(y_{0})$ $+h\sum_{i=1}^{s}b_{i}F^{\prime}(Y_{i})f(Y_{i})$
$+\frac{h^{2}}{2}\sum_{i,j=1}^{s}(b_{i}b_{j}-b_{i}a_{ij}-b_{j}a_{ji})F^{\prime\prime}\bigl{(}f(Y_{i}),f(Y_{j})\bigr{)}.$ (3b)
Proof.
Expanding $F(Y_{i})$ around $y_{0}$ using (1a) gives
$F(Y_{i})=F(y_{0})+h\sum_{j=1}^{s}a_{ij}F^{\prime}(y_{0})f(Y_{j})+\frac{h^{2}}{2}\sum_{j,k=1}^{s}a_{ij}a_{ik}F^{\prime\prime}\bigl{(}f(Y_{j}),f(Y_{k})\bigr{)}.$ (4)
For the linear terms, we can similarly expand $F^{\prime}(Y_{j})$ around $y_{0}$ to obtain
$F^{\prime}(Y_{j})=F^{\prime}(y_{0})+h\sum_{k=1}^{s}a_{jk}F^{\prime\prime}\bigl{(}-,f(Y_{k})\bigr{)}.$
Therefore,
$h\sum_{j=1}^{s}a_{ij}F^{\prime}(y_{0})f(Y_{j})$ $=h\sum_{j=1}^{s}a_{ij}F^{\prime}(Y_{j})f(Y_{j})-h^{2}\sum_{j,k=1}^{s}a_{ij}a_{jk}F^{\prime\prime}\bigl{(}f(Y_{j}),f(Y_{k})\bigr{)}$
$=h\sum_{j=1}^{s}a_{ij}F^{\prime}(Y_{j})f(Y_{j})-\frac{h^{2}}{2}\sum_{j,k=1}^{s}(a_{ij}a_{jk}+a_{ik}a_{kj})F^{\prime\prime}\bigl{(}f(Y_{j}),f(Y_{k})\bigr{)},$
where the last line symmetrizes the terms of the double sum in $j$ and $k$. Substituting this into (4) yields (3a). A similar argument is used to prove (3b). ∎
An immediate corollary of (3b) is a special case of the results in McLachlan and Stern [1]: a Runge–Kutta method is quadratic functionally equivariant if its coefficients satisfy the condition of Cooper [5] to preserve quadratic invariants. As is well known, this same coefficient condition also implies that the Runge–Kutta method is symplectic; cf. Sanz-Serna [6], Lasagni [7].
Corollary 2.
If the coefficients of a Runge–Kutta method satisfy
$b_{i}b_{j}-b_{i}a_{ij}-b_{j}a_{ji}=0,\qquad i,j=1,\ldots,s,$ (5)
then the method satisfies the quadratic functional equivariance condition (2). In particular, if $F^{\prime}(y)f(y)=0$ for all $y\in Y$, so that $F$ is invariant along $f$, then $F(y_{1})=F(y_{0})$.
We now turn our attention to the main question: which of these methods allow us to express (3) entirely in terms of $z=F(y)$ when $f$ is $F$-projectable? Recall that $f$ is $F$-projectable when there exists a vector field $g$ on $Z$ such that $F^{\prime}(y)f(y)=g\bigl{(}F(y)\bigr{)}$, i.e., $f$ and $g$ are $F$-related. For a quadratic functionally equivariant Runge–Kutta method, we may then write (3b) as
$z_{1}=z_{0}+h\sum_{i=1}^{s}b_{i}g(Z_{i}),$
where $z_{1}\coloneqq F(y_{1})$, $z_{0}\coloneqq F(y_{0})$, and $Z_{i}\coloneqq F(Y_{i})$. It remains only to consider when we can compute the internal stages $Z_{i}$ entirely on $Z$, without first computing the stages $Y_{i}$.
As an initial idea in this direction, one might think to consider methods where
$a_{ij}a_{ik}-a_{ij}a_{jk}-a_{ik}a_{kj}=0,\qquad i,j,k=1,\ldots,s,$
which would cause the quadratic terms of (3a) to vanish, just as (5) does for (3b). However, this is too much to hope for: taking $i=j=k$ shows that $a_{ii}=0$ for all $i$; taking $j=k$ then shows that $a_{ij}=0$ for all $i,j$; and taking $i=j$ in (5) finally shows that $b_{i}=0$ for all $i$. Hence, this condition is too strong to be satisfied by any nontrivial Runge–Kutta method. Instead, we consider methods satisfying the following slightly weaker condition.
Definition 3.
A Runge–Kutta method is quadratic projectable if it satisfies (5) and
$a_{ij}a_{ik}-a_{ij}a_{jk}-a_{ik}a_{kj}=0,\qquad i,j,k=1,\ldots,s\quad(j\neq k).$ (6)
For a quadratic projectable Runge–Kutta method, we may then write (3a) as
$Z_{i}$ $=z_{0}+h\sum_{j=1}^{s}a_{ij}g(Z_{j})+\frac{h^{2}}{2}\sum_{j=1}^{s}(a_{ij}^{2}-2a_{ij}a_{jj})F^{\prime\prime}\bigl{(}f(Y_{j}),f(Y_{j})\bigr{)}$
$=z_{0}+h\sum_{j=1}^{s}a_{ij}\biggl{[}g(Z_{j})+\frac{h}{2}\bigl{(}a_{ij}-2a_{jj}\bigr{)}F^{\prime\prime}\bigl{(}f(Y_{j}),f(Y_{j})\bigr{)}\biggr{]},$
since only the quadratic terms with $j=k$ remain in the sum. If we have the additional assumption (discussed further in Section 3) that $F^{\prime\prime}\bigl{(}f(y),f(y)\bigr{)}=\gamma\bigl{(}F(y)\bigr{)}$ for some vector field $\gamma$ on $Z$, then altogether
$Z_{i}=z_{0}+h\sum_{j=1}^{s}a_{ij}\biggl{[}g(Z_{j})+\frac{h}{2}\bigl{(}a_{ij}-2a_{jj}\bigr{)}\gamma(Z_{j})\biggr{]},$
and thus the method descends to one on $Z$ involving the vector fields $g$ and $\gamma$.
Lemma 4.
If (6) holds, then we have the following properties:
  1. (a)
    For all $i\neq j$, if $a_{ij}\neq 0$, then $a_{ji}=0$.
  2. (b)
    For all $i,j,k$, if $a_{ik}\neq 0$ and $a_{kj}\neq 0$, then $a_{ij}\neq 0$.
Proof.
Taking $k=i\neq j$ in (6) gives $a_{ij}a_{ji}=0$, which proves (a). Next, if $j\neq k$, then taking $a_{ij}=0$ in (6) gives $a_{ik}a_{kj}=0$, which proves the contrapositive of (b). Finally, if $j=k$, then (b) holds trivially. ∎
We are now ready to prove the main result, establishing the relationship between quadratic projectable Runge–Kutta methods and SyDIRK methods. First, we recall the definitions of explicit Runge–Kutta, DIRK, and SyDIRK methods.
Definition 5.
Suppose a Runge–Kutta method satisfies $a_{ij}=0$ whenever $j>i$. Such a method is called explicit if $a_{ii}=0$ for all $i$ and diagonally implicit otherwise. A symplectic diagonally implicit Runge–Kutta (SyDIRK) method has coefficients
$a_{ij}=\begin{cases}b_{j},&j<i,\\ b_{i}/2,&j=i,\\ 0,&j>i,\end{cases}$ (7)
for given nonzero $b_{1},\ldots,b_{s}$.
Remark 6.
Any DIRK method satisfying the symplecticity condition (5) with nonzero $b_{1},\ldots,b_{s}$ has the form (7). Furthermore, there is no loss of generality in assuming that $b_{1},\ldots,b_{s}$ are all nonzero: if $b_{i}=0$, then (5) implies $b_{j}a_{ji}=0$ for all $j$, so stage $i$ has no affect on the numerical solution and may be removed. See Theorem VI.4.4 and the remark following its proof in Hairer et al. [8].
Theorem 7.
If a Runge–Kutta method satisfies (6), then it is equivalent (up to permutation of stages) to an explicit Runge–Kutta or DIRK method. Consequently, a Runge–Kutta method is quadratic projectable if and only if it is a SyDIRK method.
Proof.
For $i\neq j$, define $j\prec i$ whenever $a_{ij}\neq 0$. Lemma 4 says that $\prec$ is a strict partial order, which may be extended to a total order on $\{1,\ldots,s\}$. Therefore, permuting the stages to sort them according to this total order gives an explicit Runge–Kutta or DIRK method. In particular, for a quadratic projectable Runge–Kutta method, the symplecticity condition (5) also holds, so the method is a SyDIRK method.
Conversely, SyDIRK coefficients (7) clearly satisfy (5). To see that they also satisfy (6), suppose without loss of generality that $k<j$. Then $a_{jk}=b_{k}$ and $a_{kj}=0$, so
$a_{ij}a_{ik}-a_{ij}a_{jk}-a_{ik}a_{kj}=a_{ij}(a_{ik}-b_{k}).$
If $i<j$, then $a_{ij}=0$, and the expression above vanishes. Otherwise, $k<j\leq i$, so $a_{ik}=b_{k}$, and again, the expression above vanishes. Hence, (6) holds, and we conclude that SyDIRK methods are quadratic projectable. ∎
Corollary 8.
Let $f$ be a vector field on $Y$, let $F\colon Y\rightarrow Z$ be quadratic, and let $g$ and $\gamma$ be vector fields on $Z$ such that $F^{\prime}(y)f(y)=g\bigl{(}F(y)\bigr{)}$ and $F^{\prime\prime}\bigl{(}f(y),f(y)\bigr{)}=\gamma\bigl{(}F(y)\bigr{)}$ for all $y\in Y$. Then an $s$-stage SyDIRK method for $\dot{y}=f(y)$ descends to the following method for $\dot{z}=g(z)$:
$Z_{i}$ $=z_{0}+h\sum_{j=1}^{i-1}b_{j}g(Z_{j})+\frac{h}{2}b_{i}g(Z_{i})-\frac{h^{2}}{8}b_{i}^{2}\gamma(Z_{i}),$ (8a)
$z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}g(Z_{i}).$ (8b)
Proof.
This follows immediately from Lemma 1 and (7). ∎
Example 9 (matrix Lie–Poisson reduction).
Let $Y=\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$ and $Z=\mathbb{R}^{n\times n}$, where $\mathbb{R}^{n\times n}$ denotes the space of real $n\times n$ matrices, and let $F(q,p)=q^{T}p$. Suppose $H\colon Y\rightarrow\mathbb{R}$ is a Hamiltonian of the form $H=\eta\circ F$, i.e., $H(q,p)=\eta(q^{T}p)$, where $\eta\colon Z\rightarrow\mathbb{R}$ is a reduced Hamiltonian. Hamilton’s equations on $Y$ can be written as
$\dot{q}$ $=\hphantom{-}\nabla_{p}H(q,p)$ $=\hphantom{-}q\nabla\eta(q^{T}p),$ (9a)
$\dot{p}$ $=-\nabla_{q}H(q,p)$ $=-p\nabla\eta(q^{T}p)^{T}.$ (9b)
We then observe that
$\frac{\mathrm{d}}{\mathrm{d}t}(q^{T}p)$ $=\dot{q}^{T}p+q^{T}\dot{p}$
$=\nabla\eta(q^{T}p)^{T}q^{T}p-q^{T}p\nabla\eta(q^{T}p)^{T}$
$=\operatorname{ad}^{\ast}_{\nabla\eta(q^{T}p)}q^{T}p.$
In other words, the Hamiltonian vector field on $Y$,
$f(q,p)=\Bigl{(}q\nabla\eta(q^{T}p),-p\nabla\eta(q^{T}p)^{T}\Bigr{)},$
is $F$-related to the reduced Hamiltonian vector field on $Z$,
$g(z)=\operatorname{ad}^{\ast}_{\nabla\eta(z)}z,$
which are the Lie–Poisson equations on $Z\cong\mathfrak{gl}(n)^{\ast}$. Observe also that
$F^{\prime\prime}\bigl{(}f(q,p),f(q,p)\bigr{)}=-2\nabla\eta(q^{T}p)^{T}q^{T}p\nabla\eta(q^{T}p)^{T}\eqqcolon\gamma(q^{T}p),$
where $\gamma(z)\coloneqq-2\nabla\eta(z)^{T}z\nabla\eta(z)^{T}$.
From Corollary 8, it follows that a SyDIRK method for (9) on $Y$ descends to the following Lie–Poisson method on $Z$:
$Z_{i}$ $=z_{0}+h\sum_{j=1}^{i-1}b_{j}\operatorname{ad}^{\ast}_{\nabla\eta(Z_{j})}Z_{j}+\frac{h}{2}b_{i}\operatorname{ad}^{\ast}_{\nabla\eta(Z_{i})}Z_{i}+\frac{h^{2}}{4}b_{i}^{2}\nabla\eta(Z_{i})^{T}Z_{i}\nabla\eta(Z_{i})^{T},$ (10a)
$z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}\operatorname{ad}^{\ast}_{\nabla\eta(Z_{i})}Z_{i}.$ (10b)
For isospectral Hamiltonian systems, this coincides with the Lie–Poisson integrators obtained via reduction of SyDIRK methods by da Silva and Lessig [2], who give an alternative expression of (10) in terms of the Cayley transform. To see the equivalence with the formulation in [2], define the additional intermediate stages
$Z^{r}_{i}\coloneqq z_{0}+h\sum_{j=1}^{i}b_{i}\operatorname{ad}^{\ast}_{\nabla\eta(Z_{j})}Z_{j},$
where $Z^{r}_{0}=z_{0}$ and $Z^{r}_{s}=z_{1}$. We may then rewrite (10) as
$Z^{r}_{i-1}$ $=\biggl{(}I-\frac{h}{2}b_{i}\nabla\eta(Z_{i})^{T}\biggr{)}Z_{i}\biggl{(}I+\frac{h}{2}b_{i}\nabla\eta(Z_{i})^{T}\biggr{)}$ $=\operatorname{dcay}^{-1}_{hb_{i}\nabla\eta(Z_{i})^{T}}Z_{i},$
$Z^{r}_{i}$ $=\biggl{(}I+\frac{h}{2}b_{i}\nabla\eta(Z_{i})^{T}\biggr{)}Z_{i}\biggl{(}I-\frac{h}{2}b_{i}\nabla\eta(Z_{i})^{T}\biggr{)}$ $=\operatorname{dcay}^{-1}_{-hb_{i}\nabla\eta(Z_{i})^{T}}Z_{i},$
which coincides (modulo slight differences in notation) with [2, Equation 3.6].
We conclude this section with a result on the numerical evolution of quadratic observables $G(z)$ by these methods, analogous to Lemma 1. This shows that (8) is affine functionally equivariant (which is clear since affine functions in $z$ are quadratic in $y$) but not quadratic functionally equivariant, in general, due to the additional $\mathcal{O}(h^{3})$ term.
Theorem 10.
Under the hypotheses of Corollary 8, if $G=G(z)$ is quadratic, then
$G(z_{1})=G(z_{0})+h\sum_{i=1}^{s}b_{i}G^{\prime}(Z_{i})g(Z_{i})+\frac{h^{3}}{8}\sum_{i=1}^{s}b_{i}^{3}G^{\prime\prime}\bigl{(}g(Z_{i}),\gamma(Z_{i})\bigr{)}.$
Proof.
As in the proof of Lemma 1, we can use (8b) to expand
$G(z_{1})=G(z_{0})+h\sum_{i=1}^{s}b_{i}G^{\prime}(z_{0})g(Z_{i})+\frac{h^{2}}{2}\sum_{i,j=1}^{s}b_{i}b_{j}G^{\prime\prime}\bigl{(}g(Z_{i}),g(Z_{j})\bigr{)}.$
Next, we use (8a) to expand $G^{\prime}(Z_{i})$ around $z_{0}$, obtaining
$G^{\prime}(Z_{i})=G^{\prime}(z_{0})+h\sum_{j=1}^{s-1}b_{j}G^{\prime\prime}\bigl{(}-,g(Z_{j})\bigr{)}+\frac{h}{2}b_{i}G^{\prime\prime}\bigl{(}-,g(Z_{i})\bigr{)}-\frac{h^{2}}{8}b_{i}^{2}G^{\prime\prime}\bigl{(}-,\gamma(Z_{i})\bigr{)}.$
Therefore,
$h\sum_{i=1}^{s}b_{i}G^{\prime}(z_{0})g(Z_{i})=h\sum_{i=1}^{s}b_{i}G^{\prime}(Z_{i})g(Z_{i})-\frac{h^{2}}{2}\sum_{i,j=1}^{s}b_{i}b_{j}G^{\prime\prime}\bigl{(}g(Z_{i}),g(Z_{j})\bigr{)}\\ +\frac{h^{3}}{8}\sum_{i=1}^{s}b_{i}^{3}G^{\prime\prime}\bigl{(}g(Z_{i}),\gamma(Z_{i})\bigr{)}.$
Substituting this into the expression for $G(z_{1})$, the double sums cancel, yielding the claimed result. ∎
3 Quadratic projectable vector fields
In this section, we study a family of vector fields that satisfies the hypotheses of Corollary 8, for which SyDIRK methods therefore descend from $Y$ to $Z$. We begin by introducing a general setting for $F$-projectability, where an additional algebraic condition involving Jordan operator algebras allows us to write $F^{\prime\prime}(y)\bigl{(}f(y),f(y)\bigr{)}=\gamma\bigl{(}F(y)\bigr{)}$. After this, we consider a different algebraic setting involving Lie algebra actions, including the important case of momentum maps in Hamiltonian mechanics, and establish similar results on projectability and descent of SyDIRK methods.
3.1 Projectability and Jordan operator algebras
We first characterize $F$-projectability with respect to an arbitrary $C^{2}$ map $F\colon Y\rightarrow Z$, with quadratic maps being a special case.
Proposition 11.
Given a $C^{2}$ map $F\colon Y\rightarrow Z$, let $f$ be a $C^{1}$ vector field on $Y$. Then $f$ is $F$-projectable if and only if, for all $y_{0}\in Y$, there exists a continuous map $B\colon Z\rightarrow L(Z,Z)$ such that for all $y\in Y$,
$F^{\prime}(y)f(y)=F^{\prime}(y_{0})f(y_{0})+B\bigl{(}F(y)\bigr{)}\bigl{(}F(y)-F(y_{0})\bigr{)}.$ (11)
Proof.
()($\Rightarrow$)( ⇒ ) Suppose $f$ is $F$-projectable, i.e., there exists a $C^{1}$ vector field $g$ on $Z$ such that $F^{\prime}(y)f(y)=g\bigl{(}F(y)\bigr{)}$ for all $y\in Y$. Now, for any $z_{0},z\in Z$, applying the fundamental theorem of calculus and chain rule along the line segment $t\mapsto(1-t)z_{0}+tz$ gives
$g(z)$ $=g(z_{0})+\int_{0}^{1}\frac{\mathrm{d}}{\mathrm{d}t}g\bigl{(}(1-t)z_{0}+tz\bigr{)}\,\mathrm{d}t$
$=g(z_{0})+\biggl{[}\int_{0}^{1}g^{\prime}\bigl{(}(1-t)z_{0}+tz\bigr{)}\,\mathrm{d}t\biggr{]}(z-z_{0});$
see Abraham et al. [9, Proposition 2.4.7]. Let $B(z)=\int_{0}^{1}g^{\prime}\bigl{(}(1-t)z_{0}+tz\bigr{)}\,\mathrm{d}t\in L(Z,Z)$, which allows the above to be written as
$g(z)=g(z_{0})+B(z)(z-z_{0}).$
(This is essentially the $C^{1}$ case of Taylor’s Theorem applied to $g$.) Taking $z_{0}=F(y_{0})$ and $z=F(y)$, then applying the $F$-relatedness of $f$ and $g$, yields (11).
()($\Leftarrow$)( ⇐ ) Conversely, if (11) holds, then define $g$ by
$g(z)=F^{\prime}(y_{0})f(y_{0})+B(z)\bigl{(}z-F(y_{0})\bigr{)}.$
Then (11) immediately implies that $F^{\prime}(y)f(y)=g\bigl{(}F(y)\bigr{)}$, i.e., $f$ is $F$-related to $g$. ∎
Assuming $f(0)=0$, applying Taylor’s Theorem as above with $y_{0}=0$ lets us write
$f(y)=A(y)y,$
where $A\colon Y\rightarrow L(Y,Y)$. If additionally $F(0)=0$, then Proposition 11 implies that an $F$-related vector field $g$ has the form
$g(z)=B(z)z.$
We now specialize to the case where $A$ factors through $F$, i.e., $A=\alpha\circ F$ for some $\alpha\colon Z\rightarrow L(Y,Y)$. In this case, we establish a sufficient condition for $F$-projectability, with additional consequences when $\alpha$ takes values in a Jordan operator algebra on $Y$.
Theorem 12.
Let $f(y)=\alpha\bigl{(}F(y)\bigr{)}y$, where $\alpha\colon Z\rightarrow\mathcal{A}\subset L(Y,Y)$.
  1. (a)
    If there exists $\beta\colon\mathcal{A}\rightarrow L(Z,Z)$ such that
    $F^{\prime}(y)Ty=\beta(T)F(y),$ (12)
    for all $y\in Y$ and $T\in\mathcal{A}$, then $f$ is $F$-related to $g(z)=\beta\bigl{(}\alpha(z)\bigr{)}z$.
  2. (b)
    Additionally, if $\mathcal{A}$ is a Jordan operator algebra on $Y$, i.e., a subspace of $L(Y,Y)$ closed under the Jordan product $S\bullet T\coloneqq\frac{1}{2}(ST+TS)$, then it follows that $F^{\prime\prime}(y)\bigl{(}f(y),f(y)\bigr{)}=\gamma\bigl{(}F(y)\bigr{)}$, where
    $\gamma(z)=\Bigl{[}\beta\bigl{(}\alpha(z)\bigr{)}^{2}-\beta\bigl{(}\alpha(z)^{2}\bigr{)}\Bigr{]}z.$
Proof.
From (12) with $T=\alpha\bigl{(}F(y)\bigr{)}$, we immediately get
$F^{\prime}(y)f(y)=F^{\prime}(y)\alpha\bigl{(}F(y)\bigr{)}y=\beta\Bigl{(}\alpha\bigl{(}F(y)\bigr{)}\Bigr{)}F(y)=g\bigl{(}F(y)\bigr{)},$
which proves 12. Next, differentiating (12) with respect to $y$ in the direction $Ty$ gives
$F^{\prime\prime}(y)(Ty,Ty)+F^{\prime}(y)T^{2}y=\beta(T)F^{\prime}(y)Ty.$
By (12), the right-hand side equals $\beta(T)^{2}F(y)$. Furthermore, if $\mathcal{A}$ is closed under the Jordan product, then in particular $T\in\mathcal{A}$ implies $T^{2}=T\bullet T\in\mathcal{A}$, so another application of (12) shows that the second term on the left-hand side equals $\beta(T^{2})F(y)$. Altogether,
$F^{\prime\prime}(y)(Ty,Ty)=\bigl{[}\beta(T)^{2}-\beta(T^{2})\bigr{]}F(y),$
and taking $T=\alpha\bigl{(}F(y)\bigr{)}$ proves (b). ∎
Corollary 13.
Let $F\colon Y\rightarrow Z$ be quadratic, and let $\alpha$ and $\beta$ be as in Theorem 12, where $\mathcal{A}$ is a Jordan operator algebra on $Y$. Then an $s$-stage SyDIRK method for $\dot{y}=\alpha\bigl{(}F(y)\bigr{)}y$ descends to the following method for $\dot{z}=\beta\bigl{(}\alpha(z)\bigr{)}z$:
$Z_{i}$ $=z_{0}+h\sum_{j=1}^{i-1}b_{j}\beta\bigl{(}\alpha(Z_{j})\bigr{)}Z_{j}+\frac{h}{2}b_{i}\beta\bigl{(}\alpha(Z_{i})\bigr{)}Z_{i}-\frac{h^{2}}{8}b_{i}^{2}\Bigl{[}\beta\bigl{(}\alpha(Z_{i})\bigr{)}^{2}-\beta\bigl{(}\alpha(Z_{i})^{2}\bigr{)}\Bigr{]}Z_{i},$
$z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}\beta\bigl{(}\alpha(Z_{i})\bigr{)}Z_{i}.$
Proof.
This follows immediately from Theorem 12 together with Corollary 8. ∎
Example 14 (matrix reduction revisited).
As in Example 9, let $Y=\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$ and $Z=\mathbb{R}^{n\times n}$ with $F(q,p)=q^{T}p$. Let $\mathcal{A}$ consist of linear operators $T_{M,N}(q,p)\coloneqq(qM,pN)$ for $M,N\in\mathbb{R}^{n\times n}$, i.e., $Y$ acting on itself by right multiplication. Then
$F^{\prime}(q,p)T_{M,N}(q,p)=(qM)^{T}p+q^{T}(pN)=M^{T}(q^{T}p)+(q^{T}p)N.$
Hence, (12) holds with $\beta(T_{M,N})z=M^{T}z+zN$, and furthermore,
$\bigl{[}\beta(T_{M,N})^{2}-\beta(T_{M,N}^{2})\bigr{]}z$ $=\bigl{[}M^{T}(M^{T}z+zN)+(M^{T}z+zN)N\bigr{]}-\bigl{[}(M^{2})^{T}z+zN^{2}\bigr{]}$
$=2M^{T}zN.$
In particular, when $N=-M^{T}$, we have
$\beta(T_{M,-M^{T}})z=M^{T}z-zM^{T}=\operatorname{ad}^{\ast}_{M}z,\qquad\bigl{[}\beta(T_{M,-M^{T}})^{2}-\beta(T_{M,-M^{T}}^{2})\bigr{]}z=-2M^{T}zM^{T}.$
This generalizes the reduction procedure in Example 9 for the special case $M=\nabla\eta(z)$, i.e., $\alpha(z)=T_{\nabla\eta(z),-\nabla\eta(z)^{T}}$.
Example 15 (Lie–Poisson integrators on $\mathbb{R}^{3}$ via the Hopf map).
McLachlan, Modin, and Verdier [10] developed collective Lie–Poisson integrators on $\mathbb{R}^{3}$ by lifting to $\mathbb{R}^{4}$ and then applying symplectic Runge–Kutta methods. We now show how this fits into our framework, and how—when the integrator on $\mathbb{R}^{4}$ is a SyDIRK method—one may integrate directly on $\mathbb{R}^{3}$ without computing the lifted trajectory on $\mathbb{R}^{4}$.
Let $Y=\mathbb{H}\cong\mathbb{R}^{4}$ and $Z=\operatorname{Im}\mathbb{H}\cong\mathbb{R}^{3}$ be the real vector spaces of quaternions and pure imaginary quaternions, respectively, and define $F\colon Y\rightarrow Z$ by $F(y)=\frac{1}{4}yky^{\ast}$. (The classical Hopf fibration map is $4F$.) Here, the star denotes quaternion conjugation, where $y=y_{0}+y_{1}i+y_{2}j+y_{3}k$ has conjugate $y^{\ast}=y_{0}-y_{1}i-y_{2}j-y_{3}k$. Letting $\mathcal{A}$ consist of left-multiplication operators $L_{x}y\coloneqq xy$, it follows that
$F^{\prime}(y)L_{x}y=\frac{1}{4}(xy)ky^{\ast}+\frac{1}{4}yk(xy)^{\ast}=xF(y)+F(y)x^{\ast}.$
Hence, (12) holds with $\beta(L_{x})z=xz+zx^{\ast}$. Furthermore,
$\bigl{[}\beta(L_{x})^{2}-\beta(L_{x}^{2})\bigr{]}z=\bigl{[}x(xz+zx^{\ast})+(xz+zx^{\ast})x^{\ast}\bigr{]}-\bigl{[}(xx)z+z(x^{\ast}x^{\ast})\bigr{]}=2xzx^{\ast}.$ (13)
Now, given a reduced Hamiltonian $\eta$ on $Z$, define $\alpha(z)=L_{-\frac{1}{2}\nabla\eta(z)}$. Then
$f(y)=\alpha\bigl{(}F(y)\bigr{)}y=-\frac{1}{2}\nabla\eta\bigl{(}F(y)\bigr{)}y,$
is $F$-related to
$g(z)=\beta\bigl{(}\alpha(z)\bigr{)}z=\frac{1}{2}\bigl{(}z\nabla\eta(z)-\nabla\eta(z)z\bigr{)}=z\times\nabla\eta(z),$
where the last expression relates the commutator on $\operatorname{Im}\mathbb{H}$ to the cross product on $\mathbb{R}^{3}$. By a similar calculation to that in [10, Appendix B], $f$ is the Hamiltonian vector field for the collective Hamiltonian $H=\eta\circ F$ on $Y$, while $g$ is the reduced Hamiltonian vector field for $\eta$ on $Z$. Moreover, from (13) with $x=-\frac{1}{2}\nabla\eta(z)$, we get
$\gamma(z)=-\frac{1}{2}\nabla\eta(z)z\nabla\eta(z)=\bigl{(}z\times\nabla\eta(z)\bigr{)}\times\nabla\eta(z)+\frac{1}{2}z\bigl{\lvert}\nabla\eta(z)\bigr{\rvert}^{2},$
again relating multiplication in $\operatorname{Im}\mathbb{H}$ to the dot and cross products on $\mathbb{R}^{3}$.
Therefore, by Corollary 13, a SyDIRK method for the collective Hamiltonian on $\mathbb{R}^{4}$ descends to the following Lie–Poisson method for the reduced Hamiltonian on $\mathbb{R}^{3}$:
$Z_{i}$ $=z_{0}+h\sum_{j=1}^{i-1}b_{j}Z_{j}\times\nabla\eta(Z_{j})+\frac{h}{2}b_{i}Z_{i}\times\nabla\eta(Z_{i})$
$\qquad-\frac{h^{2}}{8}b_{i}^{2}\Bigl{[}\bigl{(}Z_{i}\times\nabla\eta(Z_{i})\bigr{)}\times\nabla\eta(Z_{i})+\frac{1}{2}Z_{i}\bigl{\lvert}\nabla\eta(Z_{i})\bigr{\rvert}^{2}\Bigr{]},$
$z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}Z_{i}\times\nabla\eta(Z_{i}).$
Examples 14 and 15 are special cases of a more general one, where $Y$ is itself an algebra and $\mathcal{A}$ consists of left- or right-multiplication operators. The following result is contained in Albert [11, Theorem 3], but the proof is sufficiently simple that we include it here.
Lemma 16.
Given an algebra $Y$, let $L_{x}$ and $R_{x}$ denote left- and right-multiplication, respectively, by $x\in Y$. That is, $L_{x}y\coloneqq xy$ and $R_{x}y\coloneqq yx$ for $y\in Y$.
  1. (a)
    If $Y$ is left alternative, i.e., $(L_{x})^{2}=L_{x^{2}}$ for all $x\in Y$, then the left-multiplication operators $\mathcal{L}(Y)\coloneqq\{L_{x}:x\in Y\}$ form a Jordan operator algebra.
  2. (b)
    If $Y$ is right alternative, i.e., $(R_{x})^{2}=R_{x^{2}}$ for all $x\in Y$, then the right-multiplication operators $\mathcal{R}(Y)\coloneqq\{R_{x}:x\in Y\}$ form a Jordan operator algebra.
Proof.
Statements (a) and (b) are clearly equivalent, so we prove only (b), following [11]. The right-alternativity condition implies $(R_{x\pm y})^{2}=R_{(x\pm y)^{2}}$, and therefore,
$\frac{1}{2}(R_{x}R_{y}+R_{y}R_{x})=\frac{1}{4}\bigl{[}(R_{x+y})^{2}-(R_{x-y})^{2}\bigr{]}=R_{\frac{1}{4}[(x+y)^{2}-(x-y)^{2}]}=R_{\frac{1}{2}(xy+yx)}.$
Hence, $\mathcal{R}(Y)$ is closed under the Jordan product. ∎
Example 17 (octonions).
Let $Y=\mathbb{O}\cong\mathbb{R}^{8}$ be the algebra of octonions, which is (both left and right) alternative but not associative. Using the Cayley–Dickson construction, we identify $\mathbb{O}\cong\mathbb{H}\oplus\mathbb{H}$ with product $(q,p)(r,s)\coloneqq(qr-s^{\ast}p,sq+pr^{\ast})$ and conjugation $(q,p)^{\ast}\coloneqq(q^{\ast},-p)$. Now, define $F\colon Y\rightarrow Z=\mathbb{R}$ by
$F(y)=\lvert y\rvert^{2}\coloneqq yy^{\ast}.$
Letting $\mathcal{A}=\mathcal{L}(Y)$, as in Example 15, we have
$F^{\prime}(y)L_{x}y=(xy)y^{\ast}+y(xy)^{\ast}=xF(y)+F(y)x^{\ast},$
where it can be shown that alternativity of the product implies $(xy)y^{\ast}=x(yy^{\ast})$ and $y(y^{\ast}x^{\ast})=(yy^{\ast})x^{\ast}$. Hence, (12) again holds with $\beta(L_{x})z=xz+zx^{\ast}$. Furthermore, the calculation in (13) is still valid for octonions, by alternativity, so we again have
$\bigl{[}\beta(L_{x})^{2}-\beta(L_{x}^{2})\bigr{]}z=2xzx^{\ast}.$
Since $z\in\mathbb{R}$ commutes with every $x\in\mathbb{O}$, we may simply write these as
$\beta(L_{x})z=2\operatorname{Re}(x)z,\qquad\bigl{[}\beta(L_{x})^{2}-\beta(L_{x}^{2})\bigr{]}z=2\lvert x\rvert^{2}z,$
where $\operatorname{Re}(x)\coloneqq(x+x^{\ast})/2$ is the real part of $x\in\mathbb{O}$.
Finally, if we take $\alpha(z)=L_{\frac{1}{2}a(z)}$ for some $a\colon Z\rightarrow Y$, then $f(y)=\frac{1}{2}a\bigl{(}F(y)\bigr{)}y$ is $F$-related to $g(z)=\operatorname{Re}\bigl{(}a(z)\bigr{)}z$, and we have $\gamma(z)=\frac{1}{2}\bigl{\lvert}a(z)\bigr{\rvert}^{2}z$. Hence, a SyDIRK method for $\dot{y}=f(y)$ on $\mathbb{O}$ descends to the following method for $\dot{z}=g(z)$ on $\mathbb{R}$:
$Z_{i}$ $=z_{0}+h\sum_{j=1}^{i-1}b_{j}\operatorname{Re}\bigl{(}a(Z_{j})\bigr{)}Z_{j}+\frac{h}{2}b_{i}\operatorname{Re}\bigl{(}a(Z_{i})\bigr{)}Z_{i}-\frac{h^{2}}{16}b_{i}^{2}\bigl{\lvert}a(Z_{i})\bigr{\rvert}^{2}Z_{i},$
$z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}\operatorname{Re}\bigl{(}a(Z_{i})\bigr{)}Z_{i}.$
3.2 Lie algebra actions with invariant bilinear forms
We now consider quadratic observables arising from Lie algebra actions with invariant bilinear forms, for which we can prove an analog of Theorem 12 without the Jordan operator algebra assumption. This is an important setting that includes momentum maps in Hamiltonian mechanics.
Let $\mathfrak{g}$ be a Lie algebra acting linearly on $Y$. We can identify this with a subalgebra $\mathfrak{g}\subset\mathfrak{gl}(Y)$, inheriting the commutator bracket $[S,T]=ST-TS$ and acting via $y\mapsto Ty$. Suppose $\omega$ is a bilinear form on $Y$ invariant under the $\mathfrak{g}$-action, in the sense that
$\omega(Tx,y)+\omega(x,Ty)=0,$
for all $T\in\mathfrak{g}$ and $x,y\in Y$. For example, if $\omega$ is a symplectic form, then this condition means that $\mathfrak{g}$ is a subalgebra of the symplectic Lie algebra $\mathfrak{sp}(Y)$; if $\omega$ is an inner product, then $\mathfrak{g}$ is a subalgebra of the special orthogonal Lie algebra $\mathfrak{so}(Y)$.
Next, define the quadratic map $F\colon Y\rightarrow\mathfrak{g}^{\ast}$ by
$\bigl{\langle}F(y),T\bigr{\rangle}\coloneqq\frac{1}{2}\omega(Ty,y).$
This is called a momentum map of the $\mathfrak{g}$-action in the case where $\omega$ is a symplectic form on $Y$. We now obtain a projectability result for this particular form of observable.
Theorem 18.
If $\alpha\colon\mathfrak{g}^{\ast}\rightarrow\mathfrak{g}$, then $f(y)=\alpha\bigl{(}F(y)\bigr{)}y$ is $F$-related to $g(z)=\beta\bigl{(}\alpha(z)\bigr{)}$, where
$\bigl{\langle}\beta(T)z,S\bigr{\rangle}=\bigl{\langle}z,[S,T]\bigr{\rangle},$
i.e., $\beta\bigl{(}\alpha(z)\bigr{)}=-\operatorname{ad}_{\alpha(z)}^{\ast}z$. Additionally, if $\mathfrak{g}$ is closed under $(S,T)\mapsto U_{T}S\coloneqq TST$, then $F^{\prime\prime}\bigl{(}f(y),f(y)\bigr{)}=\gamma\bigl{(}F(y)\bigr{)}$, where
$\bigl{\langle}\gamma(z),S\bigr{\rangle}=-2\bigl{\langle}z,\alpha(z)S\alpha(z)\bigr{\rangle},$
i.e., $\gamma(z)=-2U_{\alpha(z)}^{\ast}z$.
Proof.
Observe that for all $S,T\in\mathfrak{g}$,
$\bigl{\langle}F^{\prime}(y)Ty,S\bigr{\rangle}$ $=\frac{1}{2}\bigl{[}\omega(STy,y)+\omega(Sy,Ty)\bigr{]}$
$=\frac{1}{2}\bigl{[}\omega(STy,y)-\omega(TSy,y)\bigr{]}$
$=\frac{1}{2}\omega\bigl{(}[S,T]y,y\bigr{)}$
$=\bigl{\langle}F(y),[S,T]\bigr{\rangle}$
$=\bigl{\langle}\beta(T)F(y),S\bigr{\rangle}.$
Hence, $F$-relatedness of $f$ and $g$ follows by taking $T=\alpha\bigl{(}F(y)\bigr{)}$. Note that this is essentially a version of the argument in Theorem 1212, where the calculation above establishes (12) with $\mathcal{A}=\mathfrak{g}$ and $Z=\mathfrak{g}^{\ast}$. Next, we have
$\bigl{\langle}F^{\prime\prime}(Ty,Ty),S\bigr{\rangle}=\omega(STy,Ty)=-\omega(TSTy,y)=-2\bigl{\langle}F(y),TST\bigr{\rangle},$
and taking $T=\alpha\bigl{(}F(y)\bigr{)}$ again completes the proof. ∎
Corollary 19.
Under the hypotheses of Theorem 18, an $s$-stage SyDIRK method for $\dot{y}=\alpha\bigl{(}F(y)\bigr{)}y$ descends to the following method for $\dot{z}=-\operatorname{ad}^{\ast}_{\alpha(z)}z$:
$Z_{i}$ $=z_{0}-h\sum_{j=1}^{i-1}b_{j}\operatorname{ad}^{\ast}_{\alpha(Z_{j})}Z_{j}-\frac{h}{2}b_{i}\operatorname{ad}^{\ast}_{\alpha(Z_{i})}Z_{i}+\frac{h^{2}}{4}b_{i}^{2}U_{\alpha(Z_{i})}^{\ast}Z_{i},$
$z_{1}$ $=z_{0}-h\sum_{i=1}^{s}b_{i}\operatorname{ad}^{\ast}_{\alpha(Z_{i})}Z_{i}.$
Remark 20.
Given any bilinear form $\omega$ on $Y$, we could define
$\mathfrak{g}=\bigl{\{}T\in\mathfrak{gl}(Y):\omega(Tx,y)+\omega(x,Ty)=0,\ \forall x,y\in Y\bigr{\}}.$
In the geometric numerical integration literature, these are sometimes referred to as quadratic Lie algebras [8, Section IV.8.3]. (Some authors only use this terminology in cases where $\omega$ is nondegenerate [12, Section 6].) In this case, it is easily seen that $S,T\in\mathfrak{g}$ implies both $[S,T]\in\mathfrak{g}$ and $U_{T}S\in\mathfrak{g}$.
Although $\mathfrak{g}$ is generally not a Jordan algebra, the quadratic representation $T\mapsto U_{T}$ also arises in the theory of quadratic Jordan algebras [13, 14]. In particular, when $\mathfrak{g}$ is defined from $\omega$ as above, then we have an associated Jordan operator algebra
$\mathcal{A}=\bigl{\{}T\in\mathfrak{gl}(Y):\omega(Tx,y)-\omega(x,Ty)=0,\ \forall x,y\in Y\bigr{\}},$
and $U_{T}S=2T\bullet(T\bullet S)-(T\bullet T)\bullet S=TST$, where again $\bullet$ is the Jordan product.
Remark 21.
Some of the earlier examples can also be viewed from the perspective of Lie algebra actions:
  • Consider the cotangent-lifted right action of $\mathfrak{g}=\mathfrak{gl}(n)$ on $Y=\mathfrak{gl}(n)\times\mathfrak{gl}(n)^{\ast}$, which we can write as $T_{M}(q,p)\coloneqq(qM,-pM^{T})$. This preserves the canonical symplectic form $\omega$ on $Y$, and the corresponding momentum map is precisely $F(q,p)=q^{T}p$, as in Examples 9 and 14, cf. Marsden and Ratiu [15, Theorem 12.1.4]. Since $\mathfrak{g}$ is a right action, note that $[T_{M_{1}},T_{M_{2}}]=-T_{[M_{1},M_{2}]}$, which explains why the expression $\operatorname{ad}^{\ast}_{M}$ on $Z=\mathbb{R}^{n\times n}$ changes signs to become $-\operatorname{ad}_{T_{M}}^{\ast}$ in Theorem 18.
  • Consider the left action of $\mathfrak{g}=\operatorname{Im}\mathbb{H}$ on $Y=\mathbb{H}$ defined by $T_{z}y\coloneqq-\frac{1}{2}zy$. This preserves the symplectic form $\omega(x,y)\coloneqq\frac{1}{2}(xky^{\ast}-ykx^{\ast})$, and the corresponding momentum map is $F(y)=\frac{1}{4}yky^{\ast}$, as in Example 15.
4 Application to matrix hydrodynamics
Euler’s equations of incompressible hydrodynamics are Lie–Poisson equations on the dual of the infinite-dimensional Lie algebra of divergence-free vector fields, cf. Arnold and Khesin [16] and references therein. For hydrodynamics in 2D (e.g., on a sphere), an approach pioneered by Zeitlin [17, 18] approximates this infinite-dimensional Lie algebra by a finite-dimensional matrix Lie algebra $\mathfrak{su}(n)$, i.e., skew-Hermitian $n\times n$ matrices with trace zero. Thus, matrix Lie–Poisson integrators on $\mathfrak{su}(n)^{\ast}$ can be used for structure-preserving simulation of 2D hydrodynamics. This is a straightforward application of Example 9 (modulo replacing $\mathbb{R}$ by $\mathbb{C}$ and transpose by adjoint) that has previously been considered by Modin and Viviani [19], da Silva and Lessig [2].
In this section, we discuss the application of the results in this paper to two generalizations of the above. The first is Lie–Poisson integration for matrix semidirect product Lie algebras, which was recently considered by Modin and Roop [20] and applied to a Zeitlin-type approximation of magnetohydrodynamics (MHD) on $\mathfrak{su}(n)\ltimes\mathfrak{su}(n)^{\ast}$. The second is an application to the 2D Navier–Stokes equations, which contain dissipation terms and are no longer Lie–Poisson (except in the inviscid case, when they coincide with Euler’s equations), and to non-conservative matrix flows more generally.
4.1 Semidirect products and magnetohydrodynamics
Modin and Roop [20] describe a model of matrix MHD given by Lie–Poisson equations on the dual of the semidirect product Lie algebra $\mathfrak{su}(n)\ltimes\mathfrak{su}(n)^{\ast}$, and they show that the Lie–Poisson integrators of Modin and Viviani [3] may be applied to this model. However, the semidirect product Lie algebra does not satisfy the assumptions of [3] to be quadratic and reductive, so the fact that the integrators remain Lie–Poisson in this case is not immediate. Instead, Modin and Roop [20] prove this by lifting to the cotangent bundle of the semidirect product Lie group.
We now show that SyDIRK methods descend to Lie–Poisson integrators for matrix semidirect product Lie algebras more generally, using a version of the argument in Examples 9 and 14 extended to complex-valued matrices.
Elements $q=(q_{1},q_{2})\in\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}$ and $p=(p_{1},p_{2})\in\bigl{(}\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}\bigr{)}^{\ast}$ may be identified with the $2n\times 2n$ triangular matrices
$q=\begin{bmatrix}q_{1}&q_{2}^{\dagger}\\ &q_{1}\end{bmatrix},\qquad p=\begin{bmatrix}p_{2}^{\dagger}\\ p_{1}&p_{2}^{\dagger}\end{bmatrix}.$
For matrices of this type, the quadratic map $F(q,p)=q^{\dagger}p$ has the form
$z=\begin{bmatrix}q_{1}^{\dagger}\\ q_{2}&q_{1}^{\dagger}\end{bmatrix}\begin{bmatrix}p_{2}^{\dagger}\\ p_{1}&p_{2}^{\dagger}\end{bmatrix}=\begin{bmatrix}(p_{2}q_{1})^{\dagger}\\ q_{1}^{\dagger}p_{1}+q_{2}p_{2}^{\dagger}&(p_{2}q_{1})^{\dagger}\end{bmatrix}\eqqcolon\begin{bmatrix}\theta^{\dagger}\\ w&\theta^{\dagger}\end{bmatrix},$
following the notation of [20] for the components $(w,\theta)\in\bigl{(}\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}\bigr{)}^{\ast}$.
Next, as in Example 14, let $T_{M,N}(q,p)\coloneqq(qM,pN)$, where
$M=\begin{bmatrix}M_{1}&M_{2}^{\dagger}\\ &M_{1}\end{bmatrix},\qquad N=\begin{bmatrix}N_{2}^{\dagger}\\ N_{1}&N_{2}^{\dagger}\end{bmatrix}.$
Then, by the same calculation as in that example, we get
$\beta(T_{M,N})z$ $=M^{\dagger}z+zN$
$=\begin{bmatrix}(\theta M_{1}+N_{2}\theta)^{\dagger}\\ M_{1}^{\dagger}w+M_{2}\theta^{\dagger}+\theta^{\dagger}N_{1}+wN_{2}^{\dagger}&(\theta M_{1}+N_{2}\theta)^{\dagger}\end{bmatrix}$
$\bigl{[}\beta(T_{M,N})^{2}-\beta(T_{M,N}^{2})\bigr{]}z$ $=2M^{\dagger}zN$
$=2\begin{bmatrix}(N_{2}\theta M_{1})^{\dagger}\\ M_{1}^{\dagger}wN_{2}^{\dagger}+M_{1}^{\dagger}\theta^{\dagger}N_{1}+M_{2}\theta^{\dagger}N_{2}^{\dagger}&(N_{2}\theta M_{1})^{\dagger}\end{bmatrix}.$
In particular, when $N=-M^{\dagger}$, we get
$\beta(T_{M,-M^{\dagger}})z$ $=[M^{\dagger},z]$
$=\begin{bmatrix}[\theta,M_{1}]^{\dagger}\\ [M_{1}^{\dagger},w]+[M_{2},\theta^{\dagger}]&[\theta,M_{1}]^{\dagger}\end{bmatrix}$
$\bigl{[}\beta(T_{M,-M^{\dagger}})^{2}-\beta(T_{M,-M^{\dagger}}^{2})\bigr{]}z$ $=-2M^{\dagger}zM^{\dagger}$
$=-2\begin{bmatrix}(M_{1}\theta M_{1})^{\dagger}\\ M_{1}^{\dagger}wM_{1}^{\dagger}+M_{1}^{\dagger}\theta^{\dagger}M_{2}+M_{2}\theta^{\dagger}M_{1}^{\dagger}&(M_{1}\theta M_{1})^{\dagger}\end{bmatrix}.$
Now, as in Modin and Roop [20], let us suppose that $\alpha(z)=T_{M(z),-M(z)^{\dagger}}$, where
$M(z)=\begin{bmatrix}M_{1}(w)&M_{2}(\theta)^{\dagger}\\ &M_{1}(w)\end{bmatrix}.$
Then it follows from Theorem 12 that
$g(w,\theta)$ $=\Bigl{(}\bigl{[}M_{1}(w)^{\dagger},w\bigr{]}+\bigl{[}M_{2}(\theta),\theta^{\dagger}\bigr{]},\bigl{[}\theta,M_{1}(w)\bigr{]}\Bigr{)}$
$\gamma(w,\theta)$ $=-2\Bigl{(}M_{1}(w)^{\dagger}wM_{1}(w)^{\dagger}+M_{1}(w)^{\dagger}\theta^{\dagger}M_{2}(\theta)+M_{2}(\theta)\theta^{\dagger}M_{1}(w)^{\dagger},M_{1}(w)\theta M_{1}(w)\Bigr{)}.$
Hence, by Corollary 13, a SyDIRK method in $y=(q,p)$ descends to the method
$W_{i}$ $=w_{0}$ $+h\sum_{j=1}^{i-1}b_{j}\Bigl{(}\bigl{[}M_{1}(W_{j})^{\dagger},W_{j}\bigr{]}+\bigl{[}M_{2}(\Theta_{j}),\Theta_{j}^{\dagger}\bigr{]}\Bigr{)}$
$+\frac{h}{2}b_{i}\Bigl{(}\bigl{[}M_{1}(W_{i})^{\dagger},W_{i}\bigr{]}+\bigl{[}M_{2}(\Theta_{i}),\Theta_{i}^{\dagger}\bigr{]}\Bigr{)}$
$+\frac{h^{2}}{4}b_{i}^{2}\Bigl{(}M_{1}(W_{i})^{\dagger}W_{i}M_{1}(W_{i})^{\dagger}+M_{1}(W_{i})^{\dagger}\Theta_{i}^{\dagger}M_{2}(\Theta_{i})+M_{2}(\Theta_{i})\Theta_{i}^{\dagger}M_{1}(W_{i})^{\dagger}\Bigr{)},$
$\Theta_{i}$ $=\theta_{0}$ $+h\sum_{j=1}^{i-1}b_{j}\bigl{[}\Theta_{j},M_{1}(W_{j})\bigr{]}+\frac{h}{2}b_{i}\bigl{[}\Theta_{i},M_{1}(W_{i})\bigr{]}+\frac{h^{2}}{4}b_{i}^{2}M_{1}(W_{i})\Theta_{i}M_{1}(W_{i}),$
$w_{1}$ $=w_{0}$ $+h\sum_{i=1}^{s}b_{i}\Bigl{(}\bigl{[}M_{1}(W_{i})^{\dagger},W_{i}\bigr{]}+\bigl{[}M_{2}(\Theta_{i}),\Theta_{i}^{\dagger}\bigr{]}\Bigr{)},$
$\theta_{1}$ $=\theta_{0}$ $+h\sum_{i=1}^{s}b_{i}\bigl{[}\Theta_{i},M_{1}(W_{i})\bigr{]}.$
Since $F$ is the momentum map of the cotangent-lifted action $M\mapsto T_{M,-M^{\dagger}}$, the fact that the SyDIRK method is symplectic implies that the descendent method is Lie–Poisson [15, Theorem 12.4.1].
In the special case when all the $n\times n$ matrix blocks are skew-Hermitian, as they are for the subalgebra $\mathfrak{su}(n)\ltimes\mathfrak{su}(n)^{\ast}$, the system $\dot{z}=g(z)$ agrees with [20, Equation 3.6]. Furthermore, the $s=1$ case of the method above (i.e., the descendent of the implicit midpoint method) coincides with the method of [20, Equations 4.6 and 4.15].
Remark 22.
In the calculation above, we have reduced along a momentum map
$\bigl{(}\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}\bigr{)}\times\bigl{(}\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}\bigr{)}^{\ast}\rightarrow\bigl{(}\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}\bigr{)}^{\ast}.$
Modin and Roop [20] reduce along a different momentum map,
$T^{\ast}\bigl{(}\mathrm{SU}(n)\ltimes\mathfrak{su}(n)^{\ast}\bigr{)}$ $\rightarrow\bigl{(}\mathfrak{su}(n)\ltimes\mathfrak{su}(n)^{\ast}\bigr{)}^{\ast}$
$(Q,m,P,\alpha)$ $\mapsto\biggl{(}\frac{PQ^{\dagger}-QP^{\dagger}}{2},Q\alpha Q^{\dagger}\biggr{)}\eqqcolon(w^{\dagger},\theta),$
corresponding to a cotangent-lifted left Lie group action. Since $m$ is irrelevant and $\alpha$ remains constant, we may restrict attention to $(Q,P)\in T^{\ast}\mathrm{SU}(n)$, where $\alpha$ is treated as a parameter of the system [20, Remark 5].
It is now possible to apply Theorem 12 and Corollary 13 in this alternative setting, where we embed $T^{\ast}\mathrm{SU}(n)$ into $Y\coloneqq\mathbb{C}^{n\times n}\times\mathbb{C}^{n\times n}$. If we define
$T_{M}(Q,P)\coloneqq(M_{1}^{\dagger}Q,M_{1}^{\dagger}P+2M_{2}^{\dagger}Q\alpha^{\dagger}),$
following [20, Equation 4.9], a calculation shows that
$\beta(T_{M})(w,\theta)$ $=(wM_{1}+M_{1}^{\dagger}w+\theta M_{2}-M_{2}^{\dagger}\theta^{\dagger},M_{1}^{\dagger}\theta+\theta M_{1}),$
$\bigl{[}\beta(T_{M})^{2}-\beta(T_{M}^{2})\bigr{]}(w,\theta)$ $=2(M_{1}^{\dagger}wM_{1}+M_{1}^{\dagger}\theta M_{2}-M_{2}^{\dagger}\theta^{\dagger}M_{1},M_{1}^{\dagger}\theta M_{1}).$
When all of these matrices are skew-Hermitian, this becomes
$\beta(T_{M})(w,\theta)$ $=\bigl{(}[M_{1}^{\dagger},w]+[M_{2},\theta^{\dagger}],[\theta,M_{1}]\bigr{)},$
$\bigl{[}\beta(T_{M})^{2}-\beta(T_{M}^{2})\bigr{]}(w,\theta)$ $=-2(M_{1}^{\dagger}wM_{1}^{\dagger}+M_{1}^{\dagger}\theta^{\dagger}M_{2}+M_{2}\theta^{\dagger}M_{1}^{\dagger},M_{1}\theta M_{1}),$
matching the previous expressions and again recovering [20, Equations 4.6 and 4.15].
4.2 General matrix flows and 2D Navier–Stokes
Suppose we wish to integrate an arbitrary matrix flow $\dot{z}=g(z)$ on $Z=\mathbb{C}^{n\times n}$, and that we would like to do so in such a way that agrees with the previously-discussed Lie–Poisson integrators when $g(z)=\bigl{[}M(z)^{\dagger},z]$. In the generic case where $z$ has distinct nonzero eigenvalues $\lambda_{1},\ldots,\lambda_{n}$, we decompose $g(z)=M(z)^{\dagger}z+zN(z)$ as follows.
We first claim that we have a direct sum decomposition
$Z=\mathcal{R}_{z}\oplus\mathcal{N}_{z},$
where $\mathcal{R}_{z}$ is the range of $L\mapsto[L,z]$ and $\mathcal{N}_{z}$ is its nullspace. By the rank-nullity theorem, it suffices to show that $\mathcal{R}_{z}\cap\mathcal{N}_{z}=\{0\}$. Note that $\mathcal{N}_{z}$ consists of matrices commuting with $z$, i.e., those that are simultaneously diagonalizable. With respect to the eigenbasis of $z$, we have $[L,z]_{ij}=(\lambda_{j}-\lambda_{i})L_{ij}$, which clearly vanishes on the diagonal $i=j$. Thus, $[L,z]$ can only be diagonal in this basis if it is identically zero, which proves the claim.
Since $z$ is invertible, multiplication by $z$ is an invertible operator on $\mathcal{N}_{z}$. It follows that any matrix can be uniquely written as $[L,z]+Pz$ for some $L\in\mathcal{R}_{z}$, $P\in\mathcal{N}_{z}$.
Remark 23.
In the eigenbasis of $z$, this decomposition is simply the splitting of a matrix into its off-diagonal and diagonal parts. Explicitly, for a matrix $K$ represented in this basis, we have
$L_{ij}=\begin{cases}0,&i=j,\\ K_{ij}/(\lambda_{j}-\lambda_{i}),&i\neq j,\end{cases}\qquad P_{ii}=\begin{cases}K_{ii}/\lambda_{i},&i=j,\\ 0,&i\neq j.\end{cases}$ (14)
Using the decomposition above, we can therefore write
$g(z)=[L,z]+Pz.$ (15)
This puts us in the situation of Example 14, where we have $F$-related vector fields
$f(q,p)=\bigl{(}qM(q^{\dagger}p),pN(q^{\dagger}p)\bigr{)},\qquad g(z)=M(z)^{\dagger}z+zN(z),$
with $F(q,p)=q^{\dagger}p$. To match (15), we take
$M(z)^{\dagger}=L+\tfrac{1}{2}P,\qquad N(z)=-L+\tfrac{1}{2}P.$ (16)
Therefore, Corollary 13 implies that a SyDIRK method for $\dot{y}=f(q,p)$ descends to the following method for $\dot{z}=g(z)$:
$Z_{i}$ $=z_{0}+h\sum_{j=1}^{i-1}b_{j}g(Z_{j})+\frac{h}{2}b_{i}g(Z_{i})-\frac{h^{2}}{4}b_{i}^{2}M(Z_{i})^{\dagger}Z_{i}N(Z_{i}),$ (17a)
$z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}g(Z_{i}).$ (17b)
Remark 24.
As with the matrix Lie–Poisson methods discussed in Example 9, we can obtain an alternative expression in terms of the additional stages
$Z^{r}_{i}\coloneqq z_{0}+h\sum_{j=1}^{i}b_{i}g(Z_{j}),$
where $Z^{r}_{0}=z_{0}$ and $Z^{r}_{s}=z_{1}$. The algorithm (17) may then be written as
$Z^{r}_{i-1}$ $=\biggl{(}I-\frac{h}{2}b_{i}M(Z_{i})^{\dagger}\biggr{)}Z_{i}\biggl{(}I-\frac{h}{2}b_{i}N(Z_{i})\biggr{)},$
$Z^{r}_{i}$ $=\biggl{(}I+\frac{h}{2}b_{i}M(Z_{i})^{\dagger}\biggr{)}Z_{i}\biggl{(}I+\frac{h}{2}b_{i}N_{i}(Z_{i})\biggr{)},$
in a similar style to the matrix Lie–Poisson methods of da Silva and Lessig [2].
To apply this approach to matrix hydrodynamics with dissipation on $\mathfrak{su}(n)$, we first consider general matrix flows on $\mathfrak{u}(n)$, the real Lie algebra of skew-Hermitian $n\times n$ matrices (with not-necessarily-vanishing trace). When $z\in\mathfrak{u}(n)$, we observe from (14) that the decomposition $g(z)=[L,z]+Pz\in\mathfrak{u}(n)$ has $L$ skew-Hermitian and $P$ Hermitian. (In fact, since the eigenbasis of $z$ is unitary, this is an orthogonal decomposition with respect to the Frobenius inner product, and we recover the “canonical splitting” of Modin and Viviani [19, 21].) It then follows from (16) that $M=N$, and therefore $\gamma(z)=2M(z)^{\dagger}zM(z)$ is skew-Hermitian as well. Hence, the method (17) remains entirely in $\mathfrak{u}(n)$. Furthermore, if $g\colon\mathfrak{u}(n)\rightarrow\mathfrak{su}(n)$, as in matrix hydrodynamics, then $z_{0}\in\mathfrak{su}(n)$ implies $z_{1}\in\mathfrak{su}(n)$, although the internal stages $Z_{i}$ may only be in $\mathfrak{u}(n)$.
Proposition 25.
The Frobenius norm of a numerical solution to (17) on $\mathfrak{u}(n)$ satisfies
$\frac{1}{2}\lVert z_{1}\rVert^{2}=\frac{1}{2}\lVert z_{0}\rVert^{2}-h\sum_{i=1}^{s}b_{i}\operatorname{tr}\Bigl{(}Z_{i}P(Z_{i})Z_{i}\Bigr{)}-\frac{h^{3}}{4}\sum_{i=1}^{s}b_{i}^{3}\operatorname{tr}\Bigl{(}M(Z_{i})^{\dagger}Z_{i}P(Z_{i})Z_{i}M(Z_{i})\Bigr{)}.$
In particular, when $g(z)=\bigl{[}M(z)^{\dagger},z\bigr{]}$, so that $P=0$, the Frobenius norm is conserved.
Proof.
We apply Theorem 10 with $G(z)=\frac{1}{2}\lVert z\rVert^{2}$. First, observe that
$G^{\prime}(z)g(z)$ $=\bigl{\langle}z,M(z)^{\dagger}z+zM(z)\bigr{\rangle}$
$=-\operatorname{tr}\Bigl{(}zM(z)^{\dagger}z+zM(z)z\Bigr{)}$
$=-\operatorname{tr}\Bigl{(}zP(z)z\Bigr{)},$
since $M^{\dagger}+M=P$. Similarly,
$G^{\prime\prime}\bigl{(}g(z),\gamma(z)\bigr{)}$ $=2\bigl{\langle}M(z)^{\dagger}z+zM(z),M(z)^{\dagger}zM(z)\bigr{\rangle}$
$=-2\operatorname{tr}\Bigl{(}M(z)^{\dagger}zM(z)^{\dagger}zM(z)+M(z)^{\dagger}zM(z)zM(z)\Bigr{)}$
$=-2\operatorname{tr}\Bigl{(}M(z)^{\dagger}zP(z)zM(z)\Bigr{)}.$
Hence, the result follows by Theorem 10. ∎
Example 26 (2D Navier–Stokes).
Let us consider the 2D Navier–Stokes equations in the vorticity formulation
$\dot{\omega}=\{\Delta^{-1}\omega,\omega\}+\nu\Delta\omega.$
Following Zeitlin [18], these equations admit a spatial semidiscretization in $\mathfrak{su}(n)$,
$\dot{w}=[\Delta_{n}^{-1}w,w]+\nu\Delta_{n}w\eqqcolon g(w).$ (18)
Here, $\Delta_{n}$ is a discrete approximation to the Laplacian (e.g., that of Hoppe and Yau [22] on the sphere), which is a self-adjoint operator on $\mathfrak{u}(n)$ but not simply an $n\times n$ matrix. The nullspace of $\Delta_{n}$ consists of imaginary multiples of the identity matrix, so it is invertible on $\mathfrak{su}(n)$. To make sense of “$\Delta_{n}^{-1}w$” for $w\in\mathfrak{u}(n)$ more generally—as we will need to do at the internal stages—we can do one of two equivalent things:
  • Take the Moore–Penrose pseudoinverse, i.e., $\Delta_{n}^{-1}\bigl{(}w-\frac{1}{n}(\operatorname{tr}w)I\bigr{)}$.
  • View $\Delta_{n}$ as an invertible operator on the quotient $\mathfrak{pu}(n)=\mathfrak{u}(n)/i\mathbb{R}I$, as in Modin and Viviani [21, Section 3], and apply $\Delta_{n}^{-1}$ to the equivalence class of $w$. The representative of this equivalence class in $\mathfrak{su}(n)$ is exactly $w-\frac{1}{n}(\operatorname{tr}w)I$, as above.
Since the Lie bracket annihilates multiples of the identity, the term $[\Delta_{n}^{-1}w,w]\in\mathfrak{su}(n)$ is thus well-defined for $w\in\mathfrak{u}(n)$, and we have $g\colon\mathfrak{u}(n)\rightarrow\mathfrak{su}(n)$ as needed.
It then follows from the preceding discussion that we can apply the method (17) to the system (18), where $L$ is skew-Hermitian, $P$ is Hermitian, and $M=N$, and this method is Lie–Poisson in the inviscid case $\nu=0$. Theorem 10 and Proposition 25 then give the following result on dissipation of energy $-\frac{1}{2}\langle\Delta_{n}^{-1}w,w\rangle$ and enstrophy $\frac{1}{2}\lVert w\rVert^{2}$.
Corollary 27.
For the method (17) applied to the Navier–Stokes–Zeitlin system (18), energy and enstrophy are dissipated monotonically up to $\mathcal{O}(h^{3})$ when $b_{i}\geq 0$ for all $i$, and enstrophy is conserved exactly in the inviscid case $\nu=0$.
Proof.
Calculating $G^{\prime}(w)g(w)$ when $G$ is energy, we get
$-\bigl{\langle}\Delta^{-1}_{n}w,[\Delta_{n}^{-1}w,w]+\nu\Delta_{n}w\bigr{\rangle}$ $=-\nu\langle\Delta^{-1}_{n}w,\Delta_{n}w\rangle$
$=-\nu\langle\Delta_{n}\Delta_{n}^{-1}w,w\rangle$
$=-\nu\bigl{\lVert}w-\tfrac{1}{n}(\operatorname{tr}w)I\bigr{\rVert}^{2}$
$\leq 0,$
where the last equality uses the fact that the composition of $\Delta_{n}$ with its pseudoinverse is the orthogonal projection onto $\mathfrak{su}(n)$. For enstrophy, we get
$\bigl{\langle}w,[\Delta_{n}^{-1}w,w]+\nu\Delta_{n}w\bigr{\rangle}=\nu\langle w,\Delta_{n}w\rangle\leq 0,$
since $\Delta_{n}$ is negative-semidefinite. Hence, dissipation of both quantities up to $\mathcal{O}(h^{3})$ follows from Theorem 10. In the inviscid case $\nu=0$, we have $P=0$, so exact numerical conservation of enstrophy follows from Proposition 25. ∎
Acknowledgments
The authors wish to thank Klas Modin for helpful conversations and generous feedback during the early stages of this project. We also would like to thank the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the program “Geometry, compatibility and structure preservation in computational differential equations,” supported by EPSRC grant number EP/R014604/1, as this paper grew out of conversations begun there. Ari Stern was supported by the National Science Foundation under Grant No. DMS-2208551.