### v2

# Quadratic projectable Runge–Kutta methods

Abstract.
Runge–Kutta methods are affine equivariant: applying a method before or after an affine change of variables yields the same numerical trajectory. However, for some applications, one would like to perform numerical integration after a quadratic change of variables. For example, in Lie–Poisson reduction, a quadratic transformation reduces the number of variables in a Hamiltonian system, yielding a more efficient representation of the dynamics. Unfortunately, directly applying a symplectic Runge–Kutta method to the reduced system generally does not preserve its Hamiltonian structure, so many proposed techniques require computing numerical trajectories of the original, unreduced system.
In this paper, we study when a Runge–Kutta method in the original variables descends to a numerical integrator expressible entirely in terms of the quadratically transformed variables. In particular, we show that symplectic diagonally implicit Runge–Kutta (SyDIRK) methods, applied to a quadratic projectable vector field, are precisely the Runge–Kutta methods that descend to a method (generally not of Runge–Kutta type) in the projected variables. We illustrate our results with several examples in both conservative and non-conservative dynamics.
2020 Mathematics Subject Classification:
Primary 37M15
1. Introduction
In computational mathematics, Runge–Kutta methods are among the most widely used classes of methods for numerical integration of ordinary differential equations. Interestingly, many of the most popular general-purpose Runge–Kutta methods have serious drawbacks when applied to the dynamics that arise in physical systems: for example, they may introduce artificial numerical dissipation into conservative systems. This has inspired a substantial line of research into structure-preserving numerical integrators that better capture the dynamics of such systems. We refer the reader to the excellent survey text by Hairer, Lubich, and Wanner [7].
Beginning with seminal work in the late 1980s, symplectic Runge–Kutta methods that preserve the structure of canonical Hamiltonian systems were characterized and found to be deeply related to those that preserve quadratic invariants [5, 22, 11, 4]. Recently, McLachlan and Stern [16] showed that such methods have important structure-preserving properties that manifest in the evolution of arbitrary quadratic observables, even those that are not invariant, with consequences for conservative and non-conservative systems alike.
The present paper is motivated by the question of when and whether it is possible to compute the numerical evolution of quadratic observables without computing a full numerical trajectory for the original system. By answering this question, we also obtain a general construction for Lie–Poisson integrators coinciding with the quadratic projection of a symplectic Runge–Kutta method onto a lower-dimensional space, allowing for more efficient computation of the corresponding Hamiltonian dynamics. We also show that such methods may be extended to non-conservative systems, including the 2D Navier–Stokes equations with dissipation arising from nonzero viscosity.
Consider a system of ordinary differential equations $\dot{y}=f(y)$ on a vector space $Y$. If $F\colon Y\rightarrow Z$ is $C^{1}$, then the chain rule implies that the observable $z=F(y)$ evolves according to $\dot{z}=F^{\prime}(y)f(y)$. The vector field $f$ is said to be $F$-projectable if there exists some vector field $g$ on $Z$ such that $F^{\prime}(y)f(y)=g\bigl(F(y)\bigr)$ for all $y\in Y$, i.e., $f$ and $g$ are $F$-related. In this case, $F$ maps solutions of $\dot{y}=f(y)$ to solutions of $\dot{z}=g(z)$.
We ask when this is true for numerical solutions when $F$ is quadratic. Specifically, suppose we integrate $\dot{y}=f(y)$ numerically using an $s$-stage Runge–Kutta method with time-step size $h$,
(1.1a) $Y_{i}$ $=y_{0}+h\sum_{j=1}^{s}a_{ij}f(Y_{j}),$
(1.1b) $y_{1}$ $=y_{0}+h\sum_{i=1}^{s}b_{i}f(Y_{i}),$
where $a_{ij}$ and $b_{i}$ are given coefficients defining the method and $Y_{i}$ are the internal stages. When does $F$ map this to a numerical solution of $\dot{z}=g(z)$ by some method on $Z$? We would like this method to be defined in terms of $z$ alone, without needing to compute a trajectory and internal stages for $y$. In that case, we say the Runge–Kutta method for $\dot{y}=f(y)$ descends to one for $\dot{z}=g(z)$.
The primary motivation for this question comes from Lie–Poisson reduction of Hamiltonian systems, arising in applications ranging from rigid body mechanics to (magneto)hydrodynamics. In this setting, $F$ is a quadratic momentum map projecting dynamics from a higher-dimensional symplectic space $Y$ onto a lower-dimensional space $Z$ equipped with a Lie–Poisson (rather than symplectic) structure. Collective symplectic integrators [14, 15] apply a symplectic Runge–Kutta method to the higher-dimensional system $\dot{y}=f(y)$, then project along $F$ to obtain a numerical solution to $\dot{z}=g(z)$ that preserves the Lie–Poisson structure. Here, our question asks: When can the projected solution be computed directly on $Z$, saving the extra computational expense associated with first computing a higher-dimensional trajectory on $Y$?
Some recent results may be understood as answering special cases of this question. da Silva and Lessig [6] proved that the isospectral symplectic Runge–Kutta (IsoSyRK) schemes of Modin and Viviani [19] descend to Lie–Poisson integrators on quadratic Lie algebras, with intermediate stages also in the Lie algebra, if the underlying Runge–Kutta scheme is a symplectic diagonally implicit Runge–Kutta (SyDIRK) method. This is in line with previous results of Viviani [24], in which the implicit midpoint method, a particular case of SyDIRK, is shown to descend for such problems.
The starting point for our approach is to observe that, if the coefficients $a_{ij}$ and $b_{i}$ are such that the Runge–Kutta method preserves quadratic invariants [5] (and thus is also symplectic [22, 11, 4]), then any quadratic observable $z=F(y)$ evolves according to
(1.2) $F(y_{1})=F(y_{0})+h\sum_{i=1}^{s}b_{i}F^{\prime}(Y_{i})f(Y_{i}),$
which McLachlan and Stern [16] call quadratic functional equivariance. Applying the $F$-relatedness of $f$ and $g$ lets us write this as
$z_{1}=z_{0}+h\sum_{i=1}^{s}b_{i}g(Z_{i}),$
where $z_{1}=F(y_{1})$, $z_{0}=F(y_{0})$, and $Z_{i}=F(Y_{i})$. It therefore remains only to determine when we can compute the internal stages $Z_{i}$ entirely on $Z$, without first computing the stages $Y_{i}$. By doing so, we not only generalize previous results on special cases of Lie–Poisson reduction, but we also obtain methods that can be applied to non-conservative systems, such as the semidiscretized 2D Navier–Stokes equations.
The paper is organized as follows:
  • Section 2 introduces quadratic projectable Runge–Kutta methods, which are symplectic $s$-stage Runge–Kutta methods satisfying additional coefficient conditions allowing them to descend to $s$-stage methods on $Z$. In addition to $F$-projectability of $f$, we also assume that there exists a vector field $\gamma$ on $Z$ satisfying $F^{\prime\prime}\bigl(f(y),f(y)\bigr)=\gamma\bigl(F(y)\bigr)$, which holds for all of our target applications.
    The main result of this section (Theorem 2.9) shows that a Runge–Kutta method is quadratic projectable if and only if it is a SyDIRK method. This immediately yields (Corollary 2.11) a novel algorithmic description of the $s$-stage methods on $Z$ to which they descend, which are generally not of Runge–Kutta type.
  • Section 3 examines the assumptions introduced in the previous section, obtaining sufficient conditions on $F$ and $f$ to determine the needed vector fields $g$ and $\gamma$ explicitly. This is done in two algebraic settings, one involving Jordan operator algebras (Theorem 3.2), the other involving Lie algebra actions with invariant bilinear forms (Theorem 3.8), with momentum maps in Hamiltonian mechanics being an important special case of the latter. In each setting, these results combine with those of the previous section to give algorithmic descriptions of the integrators to which SyDIRK methods descend (Corollaries 3.3 and 3.9 respectively), with several worked examples of each.
  • Finally, Section 4 discusses the application of these methods to computational hydrodynamics, following the matrix semidiscretization approach of Zeitlin [26, 27] where (for instance) the 2D incompressible Euler equations are approximated by the Lie–Poisson equations for the matrix Lie algebra $\mathfrak{su}(n)$ with $n$ large. While the earlier theory of IsoSyRK methods [19, 24] had already been applied to the Euler equations [18, 6], we consider two extensions that take advantage of the greater generality of the results in the present paper. The first is to Lie–Poisson integration on semidirect product matrix Lie algebras, with application to magnetohydrodynamics. The second is to arbitrary, not-necessarily-conservative matrix flows, with application to the 2D Navier–Stokes equations.
2. Quadratic projectable methods and their descendents
We begin this section with some preliminary results on the evolution of quadratic observables by Runge–Kutta methods, obtaining quadratic functional equivariance under the familiar coefficient conditions of Cooper [5], Sanz-Serna [22], and Lasagni [11]. Then, we introduce additional coefficient conditions defining quadratic projectable Runge–Kutta methods, show that these conditions are uniquely satisfied by SyDIRK methods, and describe the methods to which they descend in the projected variables.
2.1. Preliminaries: Runge–Kutta methods and quadratic observables
Let $Y$ and $Z$ be finite-dimensional [^1] We make this assumption for simplicity, and because the applications of interest are generally finite-dimensional. However, many of the results may be adapted to infinite-dimensional Banach spaces. real or complex vector spaces and $F\colon Y\rightarrow Z$ be a quadratic map, meaning that we can expand $F(y)$ around any $y_{0}\in Y$ as
$F(y)=F(y_{0})+F^{\prime}(y_{0})(y-y_{0})+\frac{1}{2}F^{\prime\prime}(y-y_{0},y-y_{0}).$
Note that we simply write $F^{\prime\prime}(v,w)$ rather than $F^{\prime\prime}(y_{0})(v,w)$, since $F^{\prime\prime}$ is a constant bilinear map for quadratic $F$, i.e., it does not depend on $y_{0}$.
The following useful lemma describes how $F(y)\in Z$ evolves numerically whenever a Runge–Kutta method is applied to $y\in Y$.
Lemma 2.1.
Suppose the Runge–Kutta method (1.1) is applied to integrate $\dot{y}=f(y)$ on $Y$. If $F\colon Y\rightarrow Z$ is quadratic, then
$F(Y_{i})$ $=F(y_{0})$ $+h\sum_{j=1}^{s}a_{ij}F^{\prime}(Y_{j})f(Y_{j})$
(2.1a) $+\frac{h^{2}}{2}\sum_{j,k=1}^{s}(a_{ij}a_{ik}-a_{ij}a_{jk}-a_{ik}a_{kj})F^{\prime\prime}\bigl(f(Y_{j}),f(Y_{k})\bigr),$
$F(y_{1})$ $=F(y_{0})$ $+h\sum_{i=1}^{s}b_{i}F^{\prime}(Y_{i})f(Y_{i})$
(2.1b) $+\frac{h^{2}}{2}\sum_{i,j=1}^{s}(b_{i}b_{j}-b_{i}a_{ij}-b_{j}a_{ji})F^{\prime\prime}\bigl(f(Y_{i}),f(Y_{j})\bigr).$
Proof.
Expanding $F(Y_{i})$ around $y_{0}$ using (1.1a) gives
(2.2) $F(Y_{i})=F(y_{0})+h\sum_{j=1}^{s}a_{ij}F^{\prime}(y_{0})f(Y_{j})+\frac{h^{2}}{2}\sum_{j,k=1}^{s}a_{ij}a_{ik}F^{\prime\prime}\bigl(f(Y_{j}),f(Y_{k})\bigr).$
For the linear terms, we can similarly expand $F^{\prime}(Y_{j})$ around $y_{0}$ to obtain
$F^{\prime}(Y_{j})v=F^{\prime}(y_{0})v+h\sum_{k=1}^{s}a_{jk}F^{\prime\prime}\bigl(v,f(Y_{k})\bigr),$
for all $v\in Y$. Therefore, taking $v=f(Y_{j})$ implies
$h\sum_{j=1}^{s}a_{ij}F^{\prime}(y_{0})f(Y_{j})\\ \begin{aligned} &=h\sum_{j=1}^{s}a_{ij}F^{\prime}(Y_{j})f(Y_{j})-h^{2}\sum_{j,k=1}^{s}a_{ij}a_{jk}F^{\prime\prime}\bigl(f(Y_{j}),f(Y_{k})\bigr)\\ &=h\sum_{j=1}^{s}a_{ij}F^{\prime}(Y_{j})f(Y_{j})-\frac{h^{2}}{2}\sum_{j,k=1}^{s}(a_{ij}a_{jk}+a_{ik}a_{kj})F^{\prime\prime}\bigl(f(Y_{j}),f(Y_{k})\bigr),\end{aligned}$
where the last line symmetrizes the terms of the double sum in $j$ and $k$. Substituting this into (2.2) yields (2.1a). A similar argument is used to prove (2.1b). ∎
An immediate corollary of (2.1b) is a special case of the results in McLachlan and Stern [16]: a Runge–Kutta method is quadratic functionally equivariant if its coefficients satisfy the condition of Cooper [5] to preserve quadratic invariants. As is well known, this same coefficient condition also implies that the Runge–Kutta method is symplectic; cf. Sanz-Serna [22], Lasagni [11].
Corollary 2.2.
If the coefficients of a Runge–Kutta method satisfy
(2.3) $b_{i}b_{j}-b_{i}a_{ij}-b_{j}a_{ji}=0,\qquad i,j=1,\ldots,s,$
then the method satisfies the quadratic functional equivariance condition (1.2). In particular, if $F^{\prime}(y)f(y)=0$ for all $y\in Y$, so that $F$ is invariant along $f$, then $F(y_{1})=F(y_{0})$.
We briefly recall the relationship between the conditions (2.3) and some familiar classes of Runge–Kutta methods. The coefficients $a_{ij}$ and $b_{i}$ of an $s$-stage Runge–Kutta method are often presented in the form of a Butcher tableau, written as
$\begin{array}[b]{|ccc}a_{11}&\cdots&a_{1s}\\ \vdots&\ddots&\vdots\\ a_{ss}&\cdots&a_{ss}\\ \hline\cr b_{1}&\cdots&b_{s}\end{array}\quad.$
If $a_{ij}=0$ whenever $j\geq i$ (i.e., the matrix of coefficients $a_{ij}$ is strictly lower-triangular), then (1.1a) gives an explicit formula for each stage $Y_{i}$ in terms of the previous stages $Y_{1},\ldots,Y_{i-1}$, and the method is said to be an explicit Runge–Kutta (ERK) method. Otherwise, (1.1a) defines a system of equations that must be solved to obtain $Y_{1},\ldots,Y_{s}$, and the method is said to be an implicit Runge–Kutta (IRK) method. An IRK method satisfying $a_{ij}=0$ whenever $j>i$ (i.e., the matrix of coefficients $a_{ij}$ has nonzero entries on the diagonal but not above it) is said to be a diagonally implicit Runge–Kutta (DIRK) method. A DIRK method allows us to solve (1.1a) for each $Y_{i}$ one-at-a-time, in sequence—similarly to an ERK method—rather than solving a larger system for all $s$ stages simultaneously, making this a fairly mild form of implicitness.
There are no ERK methods satisfying (2.3), other than the trivial method $y_{1}=y_{0}$, since taking $i=j$ in (2.3) implies $b_{i}=0$ for all $i=1,\ldots,s$. However, there are many well-known families of IRK methods satisfying (2.3), including DIRK methods.
Example 2.3 (the implicit midpoint method).
The implicit midpoint method is the $1$-stage method with coefficients $a_{11}=\frac{1}{2}$ and $b_{1}=1$, directly seen to satisfy (2.3).
Example 2.4 (Gauss–Legendre collocation methods).
The $s$-stage Gauss–Legendre collocation method, whose collocation points correspond to the roots of the degree-$s$ Legendre polynomial, is an order-$2s$ IRK method satisfying (2.3); see Hairer et al. [7, Section IV.2.1]. The $s=1$ case is the implicit midpoint method. The $s=2$ and $s=3$ methods have the corresponding Butcher tableaux
$\begin{array}[b]{|cc}\frac{1}{4}&\frac{1}{4}-\frac{\sqrt{3}}{6}\\ \frac{1}{4}+\frac{\sqrt{3}}{6}&\frac{1}{4}\\ \hline\cr\frac{1}{2}&\frac{1}{2}\end{array}\quad,\qquad\begin{array}[b]{|ccc}\frac{5}{36}&\frac{2}{9}-\frac{\sqrt{15}}{15}&\frac{5}{36}-\frac{\sqrt{15}}{30}\\ \frac{5}{36}+\frac{\sqrt{15}}{24}&\frac{2}{9}&\frac{5}{36}-\frac{\sqrt{15}}{24}\\ \frac{5}{36}+\frac{\sqrt{15}}{30}&\frac{2}{9}+\frac{\sqrt{15}}{15}&\frac{5}{36}\\ \hline\cr\frac{5}{18}&\frac{4}{9}&\frac{5}{18}\end{array}\quad.$
These methods are implicit but not diagonally implicit, except when $s=1$.
Example 2.5 (SyDIRK methods).
Given nonzero $b_{1},\ldots,b_{s}$, there is a unique diagonally implicit method, called a symplectic diagonally implicit Runge–Kutta (SyDIRK) method, satisfying (2.3), whose coefficients are given by
(2.4) $a_{ij}=\begin{cases}b_{j},&j<i,\\ b_{i}/2,&j=i,\\ 0,&j>i.\end{cases}$
The $s=1$ case is the implicit midpoint method, and it can be seen that any $s$-stage SyDIRK method is a composition of implicit midpoint steps with step sizes $hb_{1},\ldots,hb_{s}$. That is, if $\Phi_{h}$ is the map $y_{0}\mapsto y_{1}$ corresponding to the implicit midpoint method with step size $h$, then an $s$-stage SyDIRK method corresponds to the map $\Psi_{h}=\Phi_{hb_{s}}\circ\cdots\circ\Phi_{hb_{1}}$. Higher-order SyDIRK methods may therefore be obtained by applying the order theory for symmetric compositions of symmetric methods; cf. Murua and Sanz-Serna [21]. The order conditions are substantially simpler than those for arbitrary Runge–Kutta methods, which involve $a_{ij}$ and $b_{i}$. The special form (2.4) ensures that the order conditions only involve $b_{i}$, and further simplifications result from the symmetry assumption $b_{i}=b_{s-i}$.
In particular, the order-$4$ condition $\sum_{i=1}^{s}b_{i}^{3}=0$ implies that at least one of the $b_{i}$ must be negative to attain higher order than the implicit midpoint method. This is in contrast with the methods of Example 2.4, where the $b_{i}$ are the necessarily-positive Gauss–Legendre quadrature weights on $[0,1]$. Consequently, SyDIRK methods do not have the strong stability properties for stiff systems (e.g., algebraic stability) of Gauss–Legendre collocation and other symplectic methods with nonnegative $b_{i}$; (2.3) alone gives a weaker notion of stability that Cooper [5] calls “orbital stability,” which is precisely conservation of quadratic invariants. Despite this stability trade-off, higher-order SyDIRK and other symplectic composition methods remain widely used in geometric numerical integration.
The minimal-stage order-$4$ SyDIRK method with symmetric $b_{i}$ is the $s=3$ method of Yoshida [25] and Suzuki [23],
$b_{1}=b_{3}=\frac{1}{2-\sqrt[3]{2}},\qquad b_{2}=-\frac{\sqrt[3]{2}}{2-\sqrt[3]{2}},$
Interestingly, with SyDIRK and other composition methods, one can often improve performance by taking more than the minimum number of stages at a given order, allowing the coefficients to be tuned to lower the error constant. A well-known example is the order-$4$ method of Suzuki [23] with $s=5$ and coefficients
$b_{1}=b_{2}=b_{4}=b_{5}=\frac{1}{4-\sqrt[3]{4}},\qquad b_{3}=-\frac{\sqrt[3]{4}}{4-\sqrt[3]{4}}.$
See the survey in Hairer et al. [7, Section V.3.2] and references therein.
Remark 2.6.
For SyDIRK methods, there is no loss of generality in assuming that $b_{1},\ldots,b_{s}$ are all nonzero: if $b_{i}=0$, then (2.3) implies $b_{j}a_{ji}=0$ for all $j$, so stage $i$ has no affect on the numerical solution and may be removed. See Theorem VI.4.4 and the remark following its proof in Hairer et al. [7].
2.2. Quadratic projectable methods
We now turn our attention to the main question: which of these methods allow us to express (2.1) entirely in terms of $z=F(y)$ when $f$ is $F$-projectable? Recall that $f$ is $F$-projectable when there exists a vector field $g$ on $Z$ such that $F^{\prime}(y)f(y)=g\bigl(F(y)\bigr)$, i.e., $f$ and $g$ are $F$-related. For a quadratic functionally equivariant Runge–Kutta method, we may then write (2.1b) as
$z_{1}=z_{0}+h\sum_{i=1}^{s}b_{i}g(Z_{i}),$
where $z_{1}\coloneqq F(y_{1})$, $z_{0}\coloneqq F(y_{0})$, and $Z_{i}\coloneqq F(Y_{i})$. It remains only to consider when we can compute the internal stages $Z_{i}$ entirely on $Z$, without first computing the stages $Y_{i}$.
As an initial idea in this direction, one might think to consider methods where
$a_{ij}a_{ik}-a_{ij}a_{jk}-a_{ik}a_{kj}=0,\qquad i,j,k=1,\ldots,s,$
which would cause the quadratic terms of (2.1a) to vanish, just as (2.3) does for (2.1b). However, this is too much to hope for: taking $i=j=k$ shows that $a_{ii}=0$ for all $i$; taking $j=k$ then shows that $a_{ij}=0$ for all $i,j$; and taking $i=j$ in (2.3) finally shows that $b_{i}=0$ for all $i$. Hence, this condition is too strong to be satisfied by any nontrivial Runge–Kutta method. Instead, we consider methods satisfying the following slightly weaker condition.
Definition 2.7.
A Runge–Kutta method is quadratic projectable if it satisfies (2.3) and
(2.5) $a_{ij}a_{ik}-a_{ij}a_{jk}-a_{ik}a_{kj}=0,\qquad i,j,k=1,\ldots,s\quad(j\neq k).$
For a quadratic projectable Runge–Kutta method, we may then write (2.1a) as
$Z_{i}$ $=z_{0}+h\sum_{j=1}^{s}a_{ij}g(Z_{j})+\frac{h^{2}}{2}\sum_{j=1}^{s}(a_{ij}^{2}-2a_{ij}a_{jj})F^{\prime\prime}\bigl(f(Y_{j}),f(Y_{j})\bigr)$
$=z_{0}+h\sum_{j=1}^{s}a_{ij}\biggl[g(Z_{j})+\frac{h}{2}\bigl(a_{ij}-2a_{jj}\bigr)F^{\prime\prime}\bigl(f(Y_{j}),f(Y_{j})\bigr)\biggr],$
since only the quadratic terms with $j=k$ remain in the sum. If we have the additional assumption (discussed further in Section 3) that $F^{\prime\prime}\bigl(f(y),f(y)\bigr)=\gamma\bigl(F(y)\bigr)$ for some vector field $\gamma$ on $Z$, then altogether
$Z_{i}=z_{0}+h\sum_{j=1}^{s}a_{ij}\biggl[g(Z_{j})+\frac{h}{2}\bigl(a_{ij}-2a_{jj}\bigr)\gamma(Z_{j})\biggr],$
and thus the method descends to one on $Z$ involving the vector fields $g$ and $\gamma$.
We are now ready to prove the main result, establishing the relationship between quadratic projectable Runge–Kutta methods and SyDIRK methods, which we do with the aid of the following lemma.
Lemma 2.8.
If (2.5) holds, then we have the following properties:
  1. (a)
    For all $i\neq j$, if $a_{ij}\neq 0$, then $a_{ji}=0$.
  2. (b)
    For all $i,j,k$, if $a_{ik}\neq 0$ and $a_{kj}\neq 0$, then $a_{ij}\neq 0$.
Proof.
Taking $k=i\neq j$ in (2.5) gives $a_{ij}a_{ji}=0$, which proves (a). Next, if $j\neq k$, then taking $a_{ij}=0$ in (2.5) gives $a_{ik}a_{kj}=0$, which proves the contrapositive of (b). Finally, if $j=k$, then (b) holds trivially. ∎
Theorem 2.9.
If a Runge–Kutta method satisfies (2.5), then it is equivalent (up to permutation of stages) to an ERK or DIRK method. Consequently, a Runge–Kutta method is quadratic projectable if and only if it is a SyDIRK method.
Proof.
For $i\neq j$, define $j\prec i$ whenever $a_{ij}\neq 0$. Lemma 2.8 says that $\prec$ is a strict partial order, which may be extended to a total order on $\{1,\ldots,s\}$. Permuting the stages to sort them according to this total order gives a lower-triangular matrix of $a_{ij}$ coefficients, which therefore corresponds to an ERK or DIRK method. In particular, for a quadratic projectable Runge–Kutta method, the symplecticity condition (2.3) also holds, so the method is a SyDIRK method.
Conversely, SyDIRK coefficients (2.4) clearly satisfy (2.3). To see that they also satisfy (2.5), suppose without loss of generality that $k<j$. Then $a_{jk}=b_{k}$ and $a_{kj}=0$, so
$a_{ij}a_{ik}-a_{ij}a_{jk}-a_{ik}a_{kj}=a_{ij}(a_{ik}-b_{k}).$
If $i<j$, then $a_{ij}=0$, and the expression above vanishes. Otherwise, $k<j\leq i$, so $a_{ik}=b_{k}$, and again, the expression above vanishes. Hence, (2.5) holds, and we conclude that SyDIRK methods are quadratic projectable. ∎
Remark 2.10.
From the discussion of order and stability in Example 2.5, it follows that a quadratic projectable Runge–Kutta method with nonnegative $b_{i}$ has order at most $2$; higher-order methods must have at least one negative $b_{i}$.
Corollary 2.11.
Let $f$ be a vector field on $Y$, let $F\colon Y\rightarrow Z$ be quadratic, and let $g$ and $\gamma$ be vector fields on $Z$ such that $F^{\prime}(y)f(y)=g\bigl(F(y)\bigr)$ and $F^{\prime\prime}\bigl(f(y),f(y)\bigr)=\gamma\bigl(F(y)\bigr)$ for all $y\in Y$. Then an $s$-stage SyDIRK method for $\dot{y}=f(y)$ descends to the following method for $\dot{z}=g(z)$:
(2.6a) $Z_{i}$ $=z_{0}+h\sum_{j=1}^{i-1}b_{j}g(Z_{j})+\frac{h}{2}b_{i}g(Z_{i})-\frac{h^{2}}{8}b_{i}^{2}\gamma(Z_{i}),$
(2.6b) $z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}g(Z_{i}).$
Proof.
This follows immediately from Lemma 2.1 and (2.4). ∎
Example 2.12 (matrix Lie–Poisson reduction).
Let $Y=\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$ and $Z=\mathbb{R}^{n\times n}$, where $\mathbb{R}^{n\times n}$ denotes the space of real $n\times n$ matrices, and let $F(q,p)=q^{T}p$. Suppose $H\colon Y\rightarrow\mathbb{R}$ is a Hamiltonian of the form $H=\eta\circ F$, i.e., $H(q,p)=\eta(q^{T}p)$, where $\eta\colon Z\rightarrow\mathbb{R}$ is a reduced Hamiltonian. Hamilton’s equations on $Y$ can be written as
(2.7a) $\dot{q}$ $=\hphantom{-}\nabla_{p}H(q,p)$ $=\hphantom{-}q\nabla\eta(q^{T}p),$
(2.7b) $\dot{p}$ $=-\nabla_{q}H(q,p)$ $=-p\nabla\eta(q^{T}p)^{T}.$
We then observe that
$\frac{\mathrm{d}}{\mathrm{d}t}(q^{T}p)$ $=\dot{q}^{T}p+q^{T}\dot{p}$
$=\nabla\eta(q^{T}p)^{T}q^{T}p-q^{T}p\nabla\eta(q^{T}p)^{T}$
$=\operatorname{ad}^{\ast}_{\nabla\eta(q^{T}p)}q^{T}p.$
In other words, the Hamiltonian vector field on $Y$,
$f(q,p)=\Bigl(q\nabla\eta(q^{T}p),-p\nabla\eta(q^{T}p)^{T}\Bigr),$
is $F$-related to the reduced Hamiltonian vector field on $Z$,
$g(z)=\operatorname{ad}^{\ast}_{\nabla\eta(z)}z,$
which are the Lie–Poisson equations on $Z\cong\mathfrak{gl}(n)^{\ast}$. Observe also that
$F^{\prime\prime}\bigl(f(q,p),f(q,p)\bigr)=-2\nabla\eta(q^{T}p)^{T}q^{T}p\nabla\eta(q^{T}p)^{T}\eqqcolon\gamma(q^{T}p),$
where $\gamma(z)\coloneqq-2\nabla\eta(z)^{T}z\nabla\eta(z)^{T}$.
From Corollary 2.11, it follows that a SyDIRK method for (2.7) on $Y$ descends to the following Lie–Poisson method on $Z$:
(2.8a) $Z_{i}$ $=z_{0}+h\sum_{j=1}^{i-1}b_{j}\operatorname{ad}^{\ast}_{\nabla\eta(Z_{j})}Z_{j}+\frac{h}{2}b_{i}\operatorname{ad}^{\ast}_{\nabla\eta(Z_{i})}Z_{i}+\frac{h^{2}}{4}b_{i}^{2}\nabla\eta(Z_{i})^{T}Z_{i}\nabla\eta(Z_{i})^{T},$
(2.8b) $z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}\operatorname{ad}^{\ast}_{\nabla\eta(Z_{i})}Z_{i}.$
For isospectral Hamiltonian systems, this coincides with the Lie–Poisson integrators obtained via reduction of SyDIRK methods by da Silva and Lessig [6], who give an alternative expression of (2.8) in terms of the Cayley transform. To see the equivalence with the formulation in [6], define the additional intermediate stages
$Z^{r}_{i}\coloneqq z_{0}+h\sum_{j=1}^{i}b_{i}\operatorname{ad}^{\ast}_{\nabla\eta(Z_{j})}Z_{j},$
where $Z^{r}_{0}=z_{0}$ and $Z^{r}_{s}=z_{1}$. We may then rewrite (2.8) as
$Z^{r}_{i-1}$ $=\biggl(I-\frac{h}{2}b_{i}\nabla\eta(Z_{i})^{T}\biggr)Z_{i}\biggl(I+\frac{h}{2}b_{i}\nabla\eta(Z_{i})^{T}\biggr)$ $=\operatorname{dcay}^{-1}_{hb_{i}\nabla\eta(Z_{i})^{T}}Z_{i},$
$Z^{r}_{i}$ $=\biggl(I+\frac{h}{2}b_{i}\nabla\eta(Z_{i})^{T}\biggr)Z_{i}\biggl(I-\frac{h}{2}b_{i}\nabla\eta(Z_{i})^{T}\biggr)$ $=\operatorname{dcay}^{-1}_{-hb_{i}\nabla\eta(Z_{i})^{T}}Z_{i},$
which coincides (modulo slight differences in notation) with [6, Equation 3.6].
We conclude this section with a result on the numerical evolution of quadratic observables $G(z)$ by these methods, analogous to Lemma 2.1. This shows that (2.6) is affine functionally equivariant (which is clear since affine functions in $z$ are quadratic in $y$) but not quadratic functionally equivariant, in general, due to the additional $\mathcal{O}(h^{3})$ term.
Theorem 2.13.
Under the hypotheses of Corollary 2.11, if $G=G(z)$ is quadratic, then
$G(z_{1})=G(z_{0})+h\sum_{i=1}^{s}b_{i}G^{\prime}(Z_{i})g(Z_{i})+\frac{h^{3}}{8}\sum_{i=1}^{s}b_{i}^{3}G^{\prime\prime}\bigl(g(Z_{i}),\gamma(Z_{i})\bigr).$
Proof.
As in the proof of Lemma 2.1, we can use (2.6b) to expand
$G(z_{1})=G(z_{0})+h\sum_{i=1}^{s}b_{i}G^{\prime}(z_{0})g(Z_{i})+\frac{h^{2}}{2}\sum_{i,j=1}^{s}b_{i}b_{j}G^{\prime\prime}\bigl(g(Z_{i}),g(Z_{j})\bigr).$
Next, we use (2.6a) to expand $G^{\prime}(Z_{i})$ around $z_{0}$, obtaining
$G^{\prime}(Z_{i})v=G^{\prime}(z_{0})v+h\sum_{j=1}^{i-1}b_{j}G^{\prime\prime}\bigl(v,g(Z_{j})\bigr)+\frac{h}{2}b_{i}G^{\prime\prime}\bigl(v,g(Z_{i})\bigr)-\frac{h^{2}}{8}b_{i}^{2}G^{\prime\prime}\bigl(v,\gamma(Z_{i})\bigr),$
for all $v\in Z$. Therefore, taking $v=g(Z_{i})$ implies
$h\sum_{i=1}^{s}b_{i}G^{\prime}(z_{0})g(Z_{i})=h\sum_{i=1}^{s}b_{i}G^{\prime}(Z_{i})g(Z_{i})-\frac{h^{2}}{2}\sum_{i,j=1}^{s}b_{i}b_{j}G^{\prime\prime}\bigl(g(Z_{i}),g(Z_{j})\bigr)\\ +\frac{h^{3}}{8}\sum_{i=1}^{s}b_{i}^{3}G^{\prime\prime}\bigl(g(Z_{i}),\gamma(Z_{i})\bigr).$
Substituting this into the expression for $G(z_{1})$, the double sums cancel, yielding the claimed result. ∎
3. Quadratic projectable vector fields
In this section, we study a family of vector fields that satisfies the hypotheses of Corollary 2.11, for which SyDIRK methods therefore descend from $Y$ to $Z$. We begin by introducing a general setting for $F$-projectability, where an additional algebraic condition involving Jordan operator algebras allows us to write $F^{\prime\prime}(y)\bigl(f(y),f(y)\bigr)=\gamma\bigl(F(y)\bigr)$. After this, we consider a different algebraic setting involving Lie algebra actions, including the important case of momentum maps in Hamiltonian mechanics, and establish similar results on projectability and descent of SyDIRK methods.
3.1. Projectability and Jordan operator algebras
We first characterize $F$-projectability with respect to an arbitrary $C^{2}$ map $F\colon Y\rightarrow Z$, with quadratic maps being a special case. Throughout this section, we denote by $L(Y,Y)$ and $L(Z,Z)$ the spaces of linear operators on $Y$ and on $Z$, respectively.
Proposition 3.1.
Given a $C^{2}$ map $F\colon Y\rightarrow Z$, let $f$ be a $C^{1}$ vector field on $Y$. Then $f$ is $F$-projectable if and only if, for all $y_{0}\in Y$, there exists a continuous map $B\colon Z\rightarrow L(Z,Z)$ such that for all $y\in Y$,
(3.1) $F^{\prime}(y)f(y)=F^{\prime}(y_{0})f(y_{0})+B\bigl(F(y)\bigr)\bigl(F(y)-F(y_{0})\bigr).$
Proof.
$(\Rightarrow)$ Suppose $f$ is $F$-projectable, i.e., there exists a $C^{1}$ vector field $g$ on $Z$ such that $F^{\prime}(y)f(y)=g\bigl(F(y)\bigr)$ for all $y\in Y$. Now, for any $z_{0},z\in Z$, applying the fundamental theorem of calculus and chain rule along the line segment $t\mapsto(1-t)z_{0}+tz$ gives
$g(z)$ $=g(z_{0})+\int_{0}^{1}\frac{\mathrm{d}}{\mathrm{d}t}g\bigl((1-t)z_{0}+tz\bigr)\,\mathrm{d}t$
$=g(z_{0})+\biggl[\int_{0}^{1}g^{\prime}\bigl((1-t)z_{0}+tz\bigr)\,\mathrm{d}t\biggr](z-z_{0});$
see Abraham et al. [1, Proposition 2.4.7]. Let $B(z)=\int_{0}^{1}g^{\prime}\bigl((1-t)z_{0}+tz\bigr)\,\mathrm{d}t\in L(Z,Z)$, which allows the above to be written as
$g(z)=g(z_{0})+B(z)(z-z_{0}).$
(This is essentially the $C^{1}$ case of Taylor’s Theorem applied to $g$.) Taking $z_{0}=F(y_{0})$ and $z=F(y)$, then applying the $F$-relatedness of $f$ and $g$, yields (3.1).
$(\Leftarrow)$ Conversely, if (3.1) holds, then define $g$ by
$g(z)=F^{\prime}(y_{0})f(y_{0})+B(z)\bigl(z-F(y_{0})\bigr).$
Then (3.1) immediately implies that $F^{\prime}(y)f(y)=g\bigl(F(y)\bigr)$, i.e., $f$ is $F$-related to $g$. ∎
Assuming $f(0)=0$, applying Taylor’s Theorem as above with $y_{0}=0$ lets us write
$f(y)=A(y)y,$
where $A\colon Y\rightarrow L(Y,Y)$. If additionally $F(0)=0$, then Proposition 3.1 implies that an $F$-related vector field $g$ has the form
$g(z)=B(z)z.$
We now specialize to the case where $A$ factors through $F$, i.e., $A=\alpha\circ F$ for some $\alpha\colon Z\rightarrow L(Y,Y)$. In this case, we establish a sufficient condition for $F$-projectability, with additional consequences when $\alpha$ takes values in a Jordan operator algebra on $Y$.
Theorem 3.2.
Let $f(y)=\alpha\bigl(F(y)\bigr)y$, where $\alpha\colon Z\rightarrow\mathcal{A}\subset L(Y,Y)$.
  1. (a)
    If there exists $\beta\colon\mathcal{A}\rightarrow L(Z,Z)$ such that
    (3.2) $F^{\prime}(y)Ty=\beta(T)F(y),$
    for all $y\in Y$ and $T\in\mathcal{A}$, then $f$ is $F$-related to $g(z)=\beta\bigl(\alpha(z)\bigr)z$.
  2. (b)
    Additionally, if $\mathcal{A}$ is a Jordan operator algebra on $Y$, i.e., a subspace of $L(Y,Y)$ closed under the Jordan product $S\bullet T\coloneqq\frac{1}{2}(ST+TS)$, then it follows that $F^{\prime\prime}(y)\bigl(f(y),f(y)\bigr)=\gamma\bigl(F(y)\bigr)$, where
    $\gamma(z)=\Bigl[\beta\bigl(\alpha(z)\bigr)^{2}-\beta\bigl(\alpha(z)^{2}\bigr)\Bigr]z.$
Proof.
From (3.2) with $T=\alpha\bigl(F(y)\bigr)$, we immediately get
$F^{\prime}(y)f(y)=F^{\prime}(y)\alpha\bigl(F(y)\bigr)y=\beta\Bigl(\alpha\bigl(F(y)\bigr)\Bigr)F(y)=g\bigl(F(y)\bigr),$
which proves 3.2. Next, differentiating (3.2) with respect to $y$ in the direction $Ty$ gives
$F^{\prime\prime}(y)(Ty,Ty)+F^{\prime}(y)T^{2}y=\beta(T)F^{\prime}(y)Ty.$
By (3.2), the right-hand side equals $\beta(T)^{2}F(y)$. Furthermore, if $\mathcal{A}$ is closed under the Jordan product, then in particular $T\in\mathcal{A}$ implies $T^{2}=T\bullet T\in\mathcal{A}$, so another application of (3.2) shows that the second term on the left-hand side equals $\beta(T^{2})F(y)$. Altogether,
$F^{\prime\prime}(y)(Ty,Ty)=\bigl[\beta(T)^{2}-\beta(T^{2})\bigr]F(y),$
and taking $T=\alpha\bigl(F(y)\bigr)$ proves (b). ∎
Corollary 3.3.
Let $F\colon Y\rightarrow Z$ be quadratic, and let $\alpha$ and $\beta$ be as in Theorem 3.2, where $\mathcal{A}$ is a Jordan operator algebra on $Y$. Then an $s$-stage SyDIRK method for $\dot{y}=\alpha\bigl(F(y)\bigr)y$ descends to the following method for $\dot{z}=\beta\bigl(\alpha(z)\bigr)z$:
$Z_{i}$ $=z_{0}\begin{aligned} &+h\sum_{j=1}^{i-1}b_{j}\beta\bigl(\alpha(Z_{j})\bigr)Z_{j}+\frac{h}{2}b_{i}\beta\bigl(\alpha(Z_{i})\bigr)Z_{i}\\ &-\frac{h^{2}}{8}b_{i}^{2}\Bigl[\beta\bigl(\alpha(Z_{i})\bigr)^{2}-\beta\bigl(\alpha(Z_{i})^{2}\bigr)\Bigr]Z_{i},\end{aligned}$
$z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}\beta\bigl(\alpha(Z_{i})\bigr)Z_{i}.$
Proof.
This follows immediately from Theorem 3.2 together with Corollary 2.11. ∎
Example 3.4 (matrix reduction revisited).
As in Example 2.12, let $Y=\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$ and $Z=\mathbb{R}^{n\times n}$ with $F(q,p)=q^{T}p$. Let $\mathcal{A}$ consist of linear operators $T_{M,N}(q,p)\coloneqq(qM,pN)$ for $M,N\in\mathbb{R}^{n\times n}$, i.e., $Y$ acting on itself by right multiplication. Then
$F^{\prime}(q,p)T_{M,N}(q,p)=(qM)^{T}p+q^{T}(pN)=M^{T}(q^{T}p)+(q^{T}p)N.$
Hence, (3.2) holds with $\beta(T_{M,N})z=M^{T}z+zN$, and furthermore,
$\bigl[\beta(T_{M,N})^{2}-\beta(T_{M,N}^{2})\bigr]z\\ \begin{aligned} &=\bigl[M^{T}(M^{T}z+zN)+(M^{T}z+zN)N\bigr]-\bigl[(M^{2})^{T}z+zN^{2}\bigr]\\ &=2M^{T}zN.\end{aligned}$
In particular, when $N=-M^{T}$, we have
$\beta(T_{M,-M^{T}})z$ $=M^{T}z-zM^{T}=\operatorname{ad}^{\ast}_{M}z,$
$\bigl[\beta(T_{M,-M^{T}})^{2}-\beta(T_{M,-M^{T}}^{2})\bigr]z$ $=-2M^{T}zM^{T}.$
This generalizes the reduction procedure in Example 2.12 for the special case $M=\nabla\eta(z)$, i.e., $\alpha(z)=T_{\nabla\eta(z),-\nabla\eta(z)^{T}}$.
Example 3.5 (Lie–Poisson integrators on $\mathbb{R}^{3}$ via the Hopf map).
McLachlan, Modin, and Verdier [15] developed collective Lie–Poisson integrators on $\mathbb{R}^{3}$ by lifting to $\mathbb{R}^{4}$ and then applying symplectic Runge–Kutta methods. We now show how this fits into our framework, and how—when the integrator on $\mathbb{R}^{4}$ is a SyDIRK method—one may integrate directly on $\mathbb{R}^{3}$ without computing the lifted trajectory on $\mathbb{R}^{4}$.
Let $Y=\mathbb{H}\cong\mathbb{R}^{4}$ and $Z=\operatorname{Im}\mathbb{H}\cong\mathbb{R}^{3}$ be the real vector spaces of quaternions and pure imaginary quaternions, respectively, and define $F\colon Y\rightarrow Z$ by $F(y)=\frac{1}{4}yky^{\ast}$. (The classical Hopf fibration map is $4F$.) Here, the star denotes quaternion conjugation, where $y=y_{0}+y_{1}i+y_{2}j+y_{3}k$ has conjugate $y^{\ast}=y_{0}-y_{1}i-y_{2}j-y_{3}k$. Letting $\mathcal{A}$ consist of left-multiplication operators $L_{x}y\coloneqq xy$, it follows that
$F^{\prime}(y)L_{x}y=\frac{1}{4}(xy)ky^{\ast}+\frac{1}{4}yk(xy)^{\ast}=xF(y)+F(y)x^{\ast}.$
Hence, (3.2) holds with $\beta(L_{x})z=xz+zx^{\ast}$. Furthermore,
(3.3) $\bigl[\beta(L_{x})^{2}-\beta(L_{x}^{2})\bigr]z=\bigl[x(xz+zx^{\ast})+(xz+zx^{\ast})x^{\ast}\bigr]-\bigl[(xx)z+z(x^{\ast}x^{\ast})\bigr]=2xzx^{\ast}.$
Now, given a reduced Hamiltonian $\eta$ on $Z$, define $\alpha(z)=L_{-\frac{1}{2}\nabla\eta(z)}$. Then
$f(y)=\alpha\bigl(F(y)\bigr)y=-\frac{1}{2}\nabla\eta\bigl(F(y)\bigr)y,$
is $F$-related to
$g(z)=\beta\bigl(\alpha(z)\bigr)z=\frac{1}{2}\bigl(z\nabla\eta(z)-\nabla\eta(z)z\bigr)=z\times\nabla\eta(z),$
where the last expression relates the commutator on $\operatorname{Im}\mathbb{H}$ to the cross product on $\mathbb{R}^{3}$. By a similar calculation to that in [15, Appendix B], $f$ is the Hamiltonian vector field for the collective Hamiltonian $H=\eta\circ F$ on $Y$, while $g$ is the reduced Hamiltonian vector field for $\eta$ on $Z$. Moreover, from (3.3) with $x=-\frac{1}{2}\nabla\eta(z)$, we get
$\gamma(z)=-\frac{1}{2}\nabla\eta(z)z\nabla\eta(z)=\bigl(z\times\nabla\eta(z)\bigr)\times\nabla\eta(z)+\frac{1}{2}z\bigl\lvert\nabla\eta(z)\bigr\rvert^{2},$
again relating multiplication in $\operatorname{Im}\mathbb{H}$ to the dot and cross products on $\mathbb{R}^{3}$.
Therefore, by Corollary 3.3, a SyDIRK method for the collective Hamiltonian on $\mathbb{R}^{4}$ descends to the following Lie–Poisson method for the reduced Hamiltonian on $\mathbb{R}^{3}$:
$Z_{i}$ $=z_{0}\begin{aligned} &+h\sum_{j=1}^{i-1}b_{j}Z_{j}\times\nabla\eta(Z_{j})+\frac{h}{2}b_{i}Z_{i}\times\nabla\eta(Z_{i})\\ &-\frac{h^{2}}{8}b_{i}^{2}\Bigl[\bigl(Z_{i}\times\nabla\eta(Z_{i})\bigr)\times\nabla\eta(Z_{i})+\frac{1}{2}Z_{i}\bigl\lvert\nabla\eta(Z_{i})\bigr\rvert^{2}\Bigr],\end{aligned}$
$z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}Z_{i}\times\nabla\eta(Z_{i}).$
Examples 3.4 and 3.5 are special cases of a more general one, where $Y$ is itself an algebra and $\mathcal{A}$ consists of left- or right-multiplication operators. The following result is contained in Albert [2, Theorem 3], but the proof is sufficiently simple that we include it here.
Lemma 3.6.
Given an algebra $Y$, let $L_{x}$ and $R_{x}$ denote left- and right-multiplication, respectively, by $x\in Y$. That is, $L_{x}y\coloneqq xy$ and $R_{x}y\coloneqq yx$ for $y\in Y$.
  1. (a)
    If $Y$ is left alternative, i.e., $(L_{x})^{2}=L_{x^{2}}$ for all $x\in Y$, then the left-multiplication operators $\mathcal{L}(Y)\coloneqq\{L_{x}:x\in Y\}$ form a Jordan operator algebra.
  2. (b)
    If $Y$ is right alternative, i.e., $(R_{x})^{2}=R_{x^{2}}$ for all $x\in Y$, then the right-multiplication operators $\mathcal{R}(Y)\coloneqq\{R_{x}:x\in Y\}$ form a Jordan operator algebra.
Proof.
Statements (a) and (b) are clearly equivalent, so we prove only (b), following [2]. The right-alternativity condition implies $(R_{x\pm y})^{2}=R_{(x\pm y)^{2}}$, and therefore,
$\frac{1}{2}(R_{x}R_{y}+R_{y}R_{x})=\frac{1}{4}\bigl[(R_{x+y})^{2}-(R_{x-y})^{2}\bigr]=R_{\frac{1}{4}[(x+y)^{2}-(x-y)^{2}]}=R_{\frac{1}{2}(xy+yx)}.$
Hence, $\mathcal{R}(Y)$ is closed under the Jordan product. ∎
Example 3.7 (octonions).
Let $Y=\mathbb{O}\cong\mathbb{R}^{8}$ be the algebra of octonions, which is (both left and right) alternative but not associative. Using the Cayley–Dickson construction, we identify $\mathbb{O}\cong\mathbb{H}\oplus\mathbb{H}$ with product $(q,p)(r,s)\coloneqq(qr-s^{\ast}p,sq+pr^{\ast})$ and conjugation $(q,p)^{\ast}\coloneqq(q^{\ast},-p)$. Now, define $F\colon Y\rightarrow Z=\mathbb{R}$ by
$F(y)=\lvert y\rvert^{2}\coloneqq yy^{\ast}.$
Letting $\mathcal{A}=\mathcal{L}(Y)$, as in Example 3.5, we have
$F^{\prime}(y)L_{x}y=(xy)y^{\ast}+y(xy)^{\ast}=xF(y)+F(y)x^{\ast},$
where it can be shown that alternativity of the product implies $(xy)y^{\ast}=x(yy^{\ast})$ and $y(y^{\ast}x^{\ast})=(yy^{\ast})x^{\ast}$. Hence, (3.2) again holds with $\beta(L_{x})z=xz+zx^{\ast}$. Furthermore, the calculation in (3.3) is still valid for octonions, by alternativity, so we again have
$\bigl[\beta(L_{x})^{2}-\beta(L_{x}^{2})\bigr]z=2xzx^{\ast}.$
Since $z\in\mathbb{R}$ commutes with every $x\in\mathbb{O}$, we may simply write these as
$\beta(L_{x})z=2\operatorname{Re}(x)z,\qquad\bigl[\beta(L_{x})^{2}-\beta(L_{x}^{2})\bigr]z=2\lvert x\rvert^{2}z,$
where $\operatorname{Re}(x)\coloneqq(x+x^{\ast})/2$ is the real part of $x\in\mathbb{O}$.
Finally, if we take $\alpha(z)=L_{\frac{1}{2}a(z)}$ for some $a\colon Z\rightarrow Y$, then $f(y)=\frac{1}{2}a\bigl(F(y)\bigr)y$ is $F$-related to $g(z)=\operatorname{Re}\bigl(a(z)\bigr)z$, and we have $\gamma(z)=\frac{1}{2}\bigl\lvert a(z)\bigr\rvert^{2}z$. Hence, a SyDIRK method for $\dot{y}=f(y)$ on $\mathbb{O}$ descends to the following method for $\dot{z}=g(z)$ on $\mathbb{R}$:
$Z_{i}$ $=z_{0}+h\sum_{j=1}^{i-1}b_{j}\operatorname{Re}\bigl(a(Z_{j})\bigr)Z_{j}+\frac{h}{2}b_{i}\operatorname{Re}\bigl(a(Z_{i})\bigr)Z_{i}-\frac{h^{2}}{16}b_{i}^{2}\bigl\lvert a(Z_{i})\bigr\rvert^{2}Z_{i},$
$z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}\operatorname{Re}\bigl(a(Z_{i})\bigr)Z_{i}.$
3.2. Lie algebra actions with invariant bilinear forms
We now consider quadratic observables arising from Lie algebra actions with invariant bilinear forms, for which we can prove an analog of Theorem 3.2 without the Jordan operator algebra assumption. This is an important setting that includes momentum maps in Hamiltonian mechanics.
Let $\mathfrak{g}$ be a Lie algebra acting linearly on $Y$. We can identify this with a subalgebra $\mathfrak{g}\subset\mathfrak{gl}(Y)$, inheriting the commutator bracket [S,T]=STTS[S,T]=ST-TS and acting via $y\mapsto Ty$. Suppose $\omega$ is a bilinear form on $Y$ invariant under the $\mathfrak{g}$-action, in the sense that
$\omega(Tx,y)+\omega(x,Ty)=0,$
for all $T\in\mathfrak{g}$ and $x,y\in Y$. For example, if $\omega$ is a symplectic form, then this condition means that $\mathfrak{g}$ is a subalgebra of the symplectic Lie algebra $\mathfrak{sp}(Y)$; if $\omega$ is an inner product, then $\mathfrak{g}$ is a subalgebra of the special orthogonal Lie algebra $\mathfrak{so}(Y)$.
Next, define the quadratic map $F\colon Y\rightarrow\mathfrak{g}^{\ast}$ by
$\bigl\langle F(y),T\bigr\rangle\coloneqq\frac{1}{2}\omega(Ty,y).$
This is called a momentum map of the $\mathfrak{g}$-action in the case where $\omega$ is a symplectic form on $Y$. We now obtain a projectability result for this particular form of observable.
Theorem 3.8.
If $\alpha\colon\mathfrak{g}^{\ast}\rightarrow\mathfrak{g}$, then $f(y)=\alpha\bigl(F(y)\bigr)y$ is $F$-related to $g(z)=\beta\bigl(\alpha(z)\bigr)z$, where
$\bigl\langle\beta(T)z,S\bigr\rangle=\bigl\langle z,[S,T]\bigr\rangle,$
i.e., $g(z)=-\operatorname{ad}_{\alpha(z)}^{\ast}z$. Additionally, if $\mathfrak{g}$ is closed under $(S,T)\mapsto U_{T}S\coloneqq TST$, then $F^{\prime\prime}\bigl(f(y),f(y)\bigr)=\gamma\bigl(F(y)\bigr)$, where
$\bigl\langle\gamma(z),S\bigr\rangle=-2\bigl\langle z,\alpha(z)S\alpha(z)\bigr\rangle,$
i.e., $\gamma(z)=-2U_{\alpha(z)}^{\ast}z$.
Proof.
Observe that for all $S,T\in\mathfrak{g}$,
$\bigl\langle F^{\prime}(y)Ty,S\bigr\rangle$ $=\frac{1}{2}\bigl[\omega(STy,y)+\omega(Sy,Ty)\bigr]$
$=\frac{1}{2}\bigl[\omega(STy,y)-\omega(TSy,y)\bigr]$
$=\frac{1}{2}\omega\bigl([S,T]y,y\bigr)$
$=\bigl\langle F(y),[S,T]\bigr\rangle$
$=\bigl\langle\beta(T)F(y),S\bigr\rangle.$
Hence, $F$-relatedness of $f$ and $g$ follows by taking $T=\alpha\bigl(F(y)\bigr)$. Note that this is essentially a version of the argument in Theorem 3.23.2, where the calculation above establishes (3.2) with $\mathcal{A}=\mathfrak{g}$ and $Z=\mathfrak{g}^{\ast}$. Next, we have
$\bigl\langle F^{\prime\prime}(Ty,Ty),S\bigr\rangle=\omega(STy,Ty)=-\omega(TSTy,y)=-2\bigl\langle F(y),TST\bigr\rangle,$
and taking $T=\alpha\bigl(F(y)\bigr)$ again completes the proof. ∎
Corollary 3.9.
Under the hypotheses of Theorem 3.8, an $s$-stage SyDIRK method for $\dot{y}=\alpha\bigl(F(y)\bigr)y$ descends to the following method for $\dot{z}=-\operatorname{ad}^{\ast}_{\alpha(z)}z$:
$Z_{i}$ $=z_{0}-h\sum_{j=1}^{i-1}b_{j}\operatorname{ad}^{\ast}_{\alpha(Z_{j})}Z_{j}-\frac{h}{2}b_{i}\operatorname{ad}^{\ast}_{\alpha(Z_{i})}Z_{i}+\frac{h^{2}}{4}b_{i}^{2}U_{\alpha(Z_{i})}^{\ast}Z_{i},$
$z_{1}$ $=z_{0}-h\sum_{i=1}^{s}b_{i}\operatorname{ad}^{\ast}_{\alpha(Z_{i})}Z_{i}.$
Remark 3.10.
Given any bilinear form $\omega$ on $Y$, we could define
$\mathfrak{g}=\bigl\{T\in\mathfrak{gl}(Y):\omega(Tx,y)+\omega(x,Ty)=0,\ \forall x,y\in Y\bigr\}.$
In the geometric numerical integration literature, these are sometimes referred to as quadratic Lie algebras [7, Section IV.8.3]. (Some authors only use this terminology in cases where $\omega$ is nondegenerate [9, Section 6].) In this case, it is easily seen that $S,T\in\mathfrak{g}$ implies both $S,T]\in\mathfrak{g}$ and $U_{T}S\in\mathfrak{g}$.
Although $\mathfrak{g}$ is generally not a Jordan algebra, the quadratic representation $T\mapsto U_{T}$ also arises in the theory of quadratic Jordan algebras [10, 13]. In particular, when $\mathfrak{g}$ is defined from $\omega$ as above, then we have an associated Jordan operator algebra
$\mathcal{A}=\bigl\{T\in\mathfrak{gl}(Y):\omega(Tx,y)-\omega(x,Ty)=0,\ \forall x,y\in Y\bigr\},$
and $U_{T}S=2T\bullet(T\bullet S)-(T\bullet T)\bullet S=TST$, where again $\bullet$ is the Jordan product.
Remark 3.11.
Some of the earlier examples can also be viewed from the perspective of Lie algebra actions:
  • Consider the cotangent-lifted right action of $\mathfrak{g}=\mathfrak{gl}(n)$ on $Y=\mathfrak{gl}(n)\times\mathfrak{gl}(n)^{\ast}$, which we can write as $T_{M}(q,p)\coloneqq(qM,-pM^{T})$. This preserves the canonical symplectic form $\omega$ on $Y$, and the corresponding momentum map is precisely $F(q,p)=q^{T}p$, as in Examples 2.12 and 3.4, cf. Marsden and Ratiu [12, Theorem 12.1.4]. Since $\mathfrak{g}$ is a right action, note that $T_{M_{1}},T_{M_{2}}]=-T_{[M_{1},M_{2}]}$, which explains why the expression $\operatorname{ad}^{\ast}_{M}$ on $Z=\mathbb{R}^{n\times n}$ changes signs to become $\operatorname{ad}_{T_{M}}^{\ast}$ in Theorem 3.8.
  • Consider the left action of $\mathfrak{g}=\operatorname{Im}\mathbb{H}$ on $Y=\mathbb{H}$ defined by $T_{z}y\coloneqq-\frac{1}{2}zy$. This preserves the symplectic form $\omega(x,y)\coloneqq\frac{1}{2}(xky^{\ast}-ykx^{\ast})$, and the corresponding momentum map is $F(y)=\frac{1}{4}yky^{\ast}$, as in Example 3.5.
4. Application to matrix hydrodynamics
Euler’s equations of incompressible hydrodynamics are Lie–Poisson equations on the dual of the infinite-dimensional Lie algebra of divergence-free vector fields, cf. Arnold and Khesin [3] and references therein. For hydrodynamics in 2D (e.g., on a sphere), an approach pioneered by Zeitlin [26, 27] approximates this infinite-dimensional Lie algebra by a finite-dimensional matrix Lie algebra $\mathfrak{su}(n)$, i.e., skew-Hermitian $n\times n$ matrices with trace zero. Thus, matrix Lie–Poisson integrators on $\mathfrak{su}(n)^{\ast}$ can be used for structure-preserving simulation of 2D hydrodynamics. This is a straightforward application of Example 2.12 (modulo replacing $\mathbb{R}$ by $\mathbb{C}$ and transpose by adjoint) that has previously been considered by Modin and Viviani [18], da Silva and Lessig [6].
In this section, we discuss the application of the results in this paper to two generalizations of the above. The first is Lie–Poisson integration for matrix semidirect product Lie algebras, which was recently considered by Modin and Roop [17] and applied to a Zeitlin-type approximation of magnetohydrodynamics (MHD) on $\mathfrak{su}(n)\ltimes\mathfrak{su}(n)^{\ast}$. The second is an application to the 2D Navier–Stokes equations, which contain dissipation terms and are no longer Lie–Poisson (except in the inviscid case, when they coincide with Euler’s equations), and to non-conservative matrix flows more generally.
4.1. Semidirect products and magnetohydrodynamics
Modin and Roop [17] describe a model of matrix MHD given by Lie–Poisson equations on the dual of the semidirect product Lie algebra $\mathfrak{su}(n)\ltimes\mathfrak{su}(n)^{\ast}$, and they show that the Lie–Poisson integrators of Modin and Viviani [19] may be applied to this model. However, the semidirect product Lie algebra does not satisfy the assumptions of [19] to be quadratic and reductive, so the fact that the integrators remain Lie–Poisson in this case is not immediate. Instead, Modin and Roop [17] prove this by lifting to the cotangent bundle of the semidirect product Lie group.
We now show that SyDIRK methods descend to Lie–Poisson integrators for matrix semidirect product Lie algebras more generally, using a version of the argument in Examples 2.12 and 3.4 extended to complex-valued matrices.
Elements $q=(q_{1},q_{2})\in\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}$ and $p=(p_{1},p_{2})\in\bigl(\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}\bigr)^{\ast}$ may be identified with the $2n\times 2n$ triangular matrices
$q=\begin{bmatrix}q_{1}&q_{2}^{\dagger}\\ &q_{1}\end{bmatrix},\qquad p=\begin{bmatrix}p_{2}^{\dagger}\\ p_{1}&p_{2}^{\dagger}\end{bmatrix}$.
For matrices of this type, the quadratic map $F(q,p)=q^{\dagger}p$ has the form
$z=\begin{bmatrix}q_{1}^{\dagger}\\ q_{2}&q_{1}^{\dagger}\end{bmatrix}\begin{bmatrix}p_{2}^{\dagger}\\ p_{1}&p_{2}^{\dagger}\end{bmatrix}=\begin{bmatrix}(p_{2}q_{1})^{\dagger}\\ q_{1}^{\dagger}p_{1}+q_{2}p_{2}^{\dagger}&(p_{2}q_{1})^{\dagger}\end{bmatrix}\eqqcolon\begin{bmatrix}\theta^{\dagger}\\ w&\theta^{\dagger}\end{bmatrix}$,
following the notation of [17] for the components $(w,\theta)\in\bigl(\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}\bigr)^{\ast}$.
Next, as in Example 3.4, let $T_{M,N}(q,p)\coloneqq(qM,pN)$, where
$M=\begin{bmatrix}M_{1}&M_{2}^{\dagger}\\ &M_{1}\end{bmatrix},\qquad N=\begin{bmatrix}N_{2}^{\dagger}\\ N_{1}&N_{2}^{\dagger}\end{bmatrix}$.
Then, by the same calculation as in that example, we get
$\beta(T_{M,N})z$ $=M^{\dagger}z+zN$
$=\begin{bmatrix}(\theta M_{1}+N_{2}\theta)^{\dagger}\\ M_{1}^{\dagger}w+M_{2}\theta^{\dagger}+\theta^{\dagger}N_{1}+wN_{2}^{\dagger}&(\theta M_{1}+N_{2}\theta)^{\dagger}\end{bmatrix}$
$\bigl[\beta(T_{M,N})^{2}-\beta(T_{M,N}^{2})\bigr]z$ $=2M^{\dagger}zN$
$=2\begin{bmatrix}(N_{2}\theta M_{1})^{\dagger}\\ M_{1}^{\dagger}wN_{2}^{\dagger}+M_{1}^{\dagger}\theta^{\dagger}N_{1}+M_{2}\theta^{\dagger}N_{2}^{\dagger}&(N_{2}\theta M_{1})^{\dagger}\end{bmatrix}$.
In particular, when $N=-M^{\dagger}$, we get
$\beta(T_{M,-M^{\dagger}})z$ $=[M^{\dagger},z]$
$=\begin{bmatrix}[\theta,M_{1}]^{\dagger}\\ [M_{1}^{\dagger},w]+[M_{2},\theta^{\dagger}]&[\theta,M_{1}]^{\dagger}\end{bmatrix}$
$\bigl[\beta(T_{M,-M^{\dagger}})^{2}-\beta(T_{M,-M^{\dagger}}^{2})\bigr]z$ $=-2M^{\dagger}zM^{\dagger}$
$=-2\mkern-4.0mu\begin{bmatrix}(M_{1}\theta M_{1})^{\dagger}\\ M_{1}^{\dagger}wM_{1}^{\dagger}+M_{1}^{\dagger}\theta^{\dagger}M_{2}+M_{2}\theta^{\dagger}M_{1}^{\dagger}&(M_{1}\theta M_{1})^{\dagger}\end{bmatrix}\mkern-4.0mu$.
Now, as in Modin and Roop [17], let us suppose that $\alpha(z)=T_{M(z),-M(z)^{\dagger}}$, where
$M(z)=\begin{bmatrix}M_{1}(w)&M_{2}(\theta)^{\dagger}\\ &M_{1}(w)\end{bmatrix}$.
Then it follows from Theorem 3.2 that
$g(w,\theta)$ $=\Bigl(\bigl[M_{1}(w)^{\dagger},w\bigr]+\bigl[M_{2}(\theta),\theta^{\dagger}\bigr],\bigl[\theta,M_{1}(w)\bigr]\Bigr)$
$\gamma(w,\theta)$ $=-2\Bigl(\begin{aligned} &M_{1}(w)^{\dagger}wM_{1}(w)^{\dagger}+M_{1}(w)^{\dagger}\theta^{\dagger}M_{2}(\theta)+M_{2}(\theta)\theta^{\dagger}M_{1}(w)^{\dagger},\\ &M_{1}(w)\theta M_{1}(w)\Bigr).\end{aligned}$
Hence, by Corollary 3.3, a SyDIRK method in $y=(q,p)$ descends to the method
$W_{i}$ $=w_{0}$ $+h\sum_{j=1}^{i-1}b_{j}\Bigl(\bigl[M_{1}(W_{j})^{\dagger},W_{j}\bigr]+\bigl[M_{2}(\Theta_{j}),\Theta_{j}^{\dagger}\bigr]\Bigr)$
$+\frac{h}{2}b_{i}\Bigl(\bigl[M_{1}(W_{i})^{\dagger},W_{i}\bigr]+\bigl[M_{2}(\Theta_{i}),\Theta_{i}^{\dagger}\bigr]\Bigr)$
$+\frac{h^{2}}{4}b_{i}^{2}\Bigl(M_{1}(W_{i})^{\dagger}W_{i}M_{1}(W_{i})^{\dagger}\begin{aligned} &+M_{1}(W_{i})^{\dagger}\Theta_{i}^{\dagger}M_{2}(\Theta_{i})\\ &+M_{2}(\Theta_{i})\Theta_{i}^{\dagger}M_{1}(W_{i})^{\dagger}\Bigr),\end{aligned}$
$\Theta_{i}$ $=\theta_{0}$ $+h\sum_{j=1}^{i-1}b_{j}\bigl[\Theta_{j},M_{1}(W_{j})\bigr]+\frac{h}{2}b_{i}\bigl[\Theta_{i},M_{1}(W_{i})\bigr]+\frac{h^{2}}{4}b_{i}^{2}M_{1}(W_{i})\Theta_{i}M_{1}(W_{i}),$
$w_{1}$ $=w_{0}$ $+h\sum_{i=1}^{s}b_{i}\Bigl(\bigl[M_{1}(W_{i})^{\dagger},W_{i}\bigr]+\bigl[M_{2}(\Theta_{i}),\Theta_{i}^{\dagger}\bigr]\Bigr),$
$\theta_{1}$ $=\theta_{0}$ $+h\sum_{i=1}^{s}b_{i}\bigl[\Theta_{i},M_{1}(W_{i})\bigr].$
Since $F$ is the momentum map of the cotangent-lifted action $M\mapsto T_{M,-M^{\dagger}}$, the fact that the SyDIRK method is symplectic implies that the descendent method is Lie–Poisson [12, Theorem 12.4.1].
In the special case when all the $n\times n$ matrix blocks are skew-Hermitian, as they are for the subalgebra $\mathfrak{su}(n)\ltimes\mathfrak{su}(n)^{\ast}$, the system $\dot{z}=g(z)$ agrees with [17, Equation 3.8]. Furthermore, the $s=1$ case of the method above (i.e., the descendent of the implicit midpoint method) coincides with the method of [17, Equations 4.6 and 4.15].
Remark 4.1.
In the calculation above, we have reduced along a momentum map
$\bigl(\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}\bigr)\times\bigl(\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}\bigr)^{\ast}\rightarrow\bigl(\mathfrak{gl}(n,\mathbb{C})\ltimes\mathfrak{gl}(n,\mathbb{C})^{\ast}\bigr)^{\ast}.$
Modin and Roop [17] reduce along a different momentum map,
$T^{\ast}\bigl(\mathrm{SU}(n)\ltimes\mathfrak{su}(n)^{\ast}\bigr)$ $\rightarrow\bigl(\mathfrak{su}(n)\ltimes\mathfrak{su}(n)^{\ast}\bigr)^{\ast}$
$(Q,m,P,\alpha)$ $\mapsto\biggl(\frac{PQ^{\dagger}-QP^{\dagger}}{2},Q\alpha Q^{\dagger}\biggr)\eqqcolon(w^{\dagger},\theta),$
corresponding to a cotangent-lifted left Lie group action. Since $m$ is irrelevant and $\alpha$ remains constant, we may restrict attention to $(Q,P)\in T^{\ast}\mathrm{SU}(n)$, where $\alpha$ is treated as a parameter of the system [17, Remarks 7 and 8].
It is now possible to apply Theorem 3.2 and Corollary 3.3 in this alternative setting, where we embed $T^{\ast}\mathrm{SU}(n)$ into $Y\coloneqq\mathbb{C}^{n\times n}\times\mathbb{C}^{n\times n}$. If we define
$T_{M}(Q,P)\coloneqq(M_{1}^{\dagger}Q,M_{1}^{\dagger}P+2M_{2}^{\dagger}Q\alpha^{\dagger}),$
following [17, Equation 4.9], a calculation shows that
$\beta(T_{M})(w,\theta)$ $=(wM_{1}+M_{1}^{\dagger}w+\theta M_{2}-M_{2}^{\dagger}\theta^{\dagger},M_{1}^{\dagger}\theta+\theta M_{1}),$
$\bigl[\beta(T_{M})^{2}-\beta(T_{M}^{2})\bigr](w,\theta)$ $=2(M_{1}^{\dagger}wM_{1}+M_{1}^{\dagger}\theta M_{2}-M_{2}^{\dagger}\theta^{\dagger}M_{1},M_{1}^{\dagger}\theta M_{1}).$
When all of these matrices are skew-Hermitian, this becomes
$\beta(T_{M})(w,\theta)$ $=\bigl([M_{1}^{\dagger},w]+[M_{2},\theta^{\dagger}],[\theta,M_{1}]\bigr),$
$\bigl[\beta(T_{M})^{2}-\beta(T_{M}^{2})\bigr](w,\theta)$ $=-2(M_{1}^{\dagger}wM_{1}^{\dagger}+M_{1}^{\dagger}\theta^{\dagger}M_{2}+M_{2}\theta^{\dagger}M_{1}^{\dagger},M_{1}\theta M_{1}),$
matching the previous expressions and again recovering [17, Equations 4.6 and 4.15].
4.2. General matrix flows and 2D Navier–Stokes
Suppose we wish to integrate an arbitrary matrix flow $\dot{z}=g(z)$ on $Z=\mathbb{C}^{n\times n}$, and that we would like to do so in such a way that agrees with the previously-discussed Lie–Poisson integrators when $g(z)=\bigl[M(z)^{\dagger},z]$. In the generic case where $z$ has distinct nonzero eigenvalues $\lambda_{1},\ldots,\lambda_{n}$, we decompose $g(z)=M(z)^{\dagger}z+zN(z)$ as follows.
We first claim that we have a direct sum decomposition
$Z=\mathcal{R}_{z}\oplus\mathcal{N}_{z},$
where $\mathcal{R}_{z}$ is the range of $L\mapsto[L,z]$ and $\mathcal{N}_{z}$ is its nullspace. By the rank-nullity theorem, it suffices to show that $\mathcal{R}_{z}\cap\mathcal{N}_{z}=\{0\}$. Note that $\mathcal{N}_{z}$ consists of matrices commuting with $z$, i.e., those that are simultaneously diagonalizable. With respect to the eigenbasis of $z$, we have $L,z]_{ij}=(\lambda_{j}-\lambda_{i})L_{ij}$, which clearly vanishes on the diagonal $i=j$. Thus, $[L,z]$ can only be diagonal in this basis if it is identically zero, which proves the claim.
Since $z$ is invertible, multiplication by $z$ is an invertible operator on $\mathcal{N}_{z}$. It follows that any matrix can be uniquely written as $[L,z]+Pz$ for some $L\in\mathcal{R}_{z}$, $P\in\mathcal{N}_{z}$.
Remark 4.2.
In the eigenbasis of $z$, this decomposition is simply the splitting of a matrix into its off-diagonal and diagonal parts. Explicitly, for a matrix $K$ represented in this basis, we have
(4.1) $L_{ij}=\begin{cases}0,&i=j,\\ K_{ij}/(\lambda_{j}-\lambda_{i}),&i\neq j,\end{cases}\qquad P_{ii}=\begin{cases}K_{ii}/\lambda_{i},&i=j,\\ 0,&i\neq j.\end{cases}$
Using the decomposition above, we can therefore write
(4.2) $g(z)=[L,z]+Pz.$
This puts us in the situation of Example 3.4, where we have $F$-related vector fields
$f(q,p)=\bigl(qM(q^{\dagger}p),pN(q^{\dagger}p)\bigr),\qquad g(z)=M(z)^{\dagger}z+zN(z),$
with $F(q,p)=q^{\dagger}p$. To match (4.2), we take
(4.3) $M(z)^{\dagger}=L+\tfrac{1}{2}P,\qquad N(z)=-L+\tfrac{1}{2}P.$
Therefore, Corollary 3.3 implies that a SyDIRK method for $\dot{y}=f(q,p)$ descends to the following method for $\dot{z}=g(z)$:
(4.4a) $Z_{i}$ $=z_{0}+h\sum_{j=1}^{i-1}b_{j}g(Z_{j})+\frac{h}{2}b_{i}g(Z_{i})-\frac{h^{2}}{4}b_{i}^{2}M(Z_{i})^{\dagger}Z_{i}N(Z_{i}),$
(4.4b) $z_{1}$ $=z_{0}+h\sum_{i=1}^{s}b_{i}g(Z_{i}).$
Remark 4.3.
As with the matrix Lie–Poisson methods discussed in Example 2.12, we can obtain an alternative expression in terms of the additional stages
$Z^{r}_{i}\coloneqq z_{0}+h\sum_{j=1}^{i}b_{i}g(Z_{j}),$
where $Z^{r}_{0}=z_{0}$ and $Z^{r}_{s}=z_{1}$. The algorithm (4.4) may then be written as
$Z^{r}_{i-1}$ $=\biggl(I-\frac{h}{2}b_{i}M(Z_{i})^{\dagger}\biggr)Z_{i}\biggl(I-\frac{h}{2}b_{i}N(Z_{i})\biggr),$
$Z^{r}_{i}$ $=\biggl(I+\frac{h}{2}b_{i}M(Z_{i})^{\dagger}\biggr)Z_{i}\biggl(I+\frac{h}{2}b_{i}N(Z_{i})\biggr),$
in a similar style to the matrix Lie–Poisson methods of da Silva and Lessig [6].
To apply this approach to matrix hydrodynamics with dissipation on $\mathfrak{su}(n)$, we first consider general matrix flows on $\mathfrak{u}(n)$, the real Lie algebra of skew-Hermitian $n\times n$ matrices (with not-necessarily-vanishing trace). When $z\in\mathfrak{u}(n)$, we observe from (4.1) that the decomposition $g(z)=[L,z]+Pz\in\mathfrak{u}(n)$ has $L$ skew-Hermitian and $P$ Hermitian. (In fact, since the eigenbasis of $z$ is unitary, this is an orthogonal decomposition with respect to the Frobenius inner product, and we recover the “canonical splitting” of Modin and Viviani [18, 20].) It then follows from (4.3) that $M=N$, and therefore $\gamma(z)=2M(z)^{\dagger}zM(z)$ is skew-Hermitian as well. Hence, the method (4.4) remains entirely in $\mathfrak{u}(n)$. Furthermore, if $g\colon\mathfrak{u}(n)\rightarrow\mathfrak{su}(n)$, as in matrix hydrodynamics, then $z_{0}\in\mathfrak{su}(n)$ implies $z_{1}\in\mathfrak{su}(n)$, although the internal stages $Z_{i}$ may only be in $\mathfrak{u}(n)$.
Proposition 4.4.
The Frobenius norm of a numerical solution to (4.4) on $\mathfrak{u}(n)$ satisfies
$\frac{1}{2}\lVert z_{1}\rVert^{2}=\frac{1}{2}\lVert z_{0}\rVert^{2}-h\sum_{i=1}^{s}b_{i}\operatorname{tr}\Bigl(Z_{i}P(Z_{i})Z_{i}\Bigr)-\frac{h^{3}}{4}\sum_{i=1}^{s}b_{i}^{3}\operatorname{tr}\Bigl(M(Z_{i})^{\dagger}Z_{i}P(Z_{i})Z_{i}M(Z_{i})\Bigr).$
In particular, when $g(z)=\bigl[M(z)^{\dagger},z\bigr]$, so that $P=0$, the Frobenius norm is conserved.
Proof.
We apply Theorem 2.13 with $G(z)=\frac{1}{2}\lVert z\rVert^{2}$. First, observe that
$G^{\prime}(z)g(z)$ $=\bigl\langle z,M(z)^{\dagger}z+zM(z)\bigr\rangle$
$=-\operatorname{tr}\Bigl(zM(z)^{\dagger}z+zM(z)z\Bigr)$
$=-\operatorname{tr}\Bigl(zP(z)z\Bigr),$
since $M^{\dagger}+M=P$. Similarly,
$G^{\prime\prime}\bigl(g(z),\gamma(z)\bigr)$ $=2\bigl\langle M(z)^{\dagger}z+zM(z),M(z)^{\dagger}zM(z)\bigr\rangle$
$=-2\operatorname{tr}\Bigl(M(z)^{\dagger}zM(z)^{\dagger}zM(z)+M(z)^{\dagger}zM(z)zM(z)\Bigr)$
$=-2\operatorname{tr}\Bigl(M(z)^{\dagger}zP(z)zM(z)\Bigr).$
Hence, the result follows by Theorem 2.13. ∎
Example 4.5 (2D Navier–Stokes).
Let us consider the 2D Navier–Stokes equations in the vorticity formulation
$\dot{\omega}=\{\Delta^{-1}\omega,\omega\}+\nu\Delta\omega.$
Following Zeitlin [27], these equations admit a spatial semidiscretization in $\mathfrak{su}(n)$,
(4.5) $\dot{w}=[\Delta_{n}^{-1}w,w]+\nu\Delta_{n}w\eqqcolon g(w).$
Here, $\Delta_{n}$ is a discrete approximation to the Laplacian (e.g., that of Hoppe and Yau [8] on the sphere), which is a self-adjoint operator on $\mathfrak{u}(n)$ but not simply an $n\times n$ matrix. The nullspace of $\Delta_{n}$ consists of imaginary multiples of the identity matrix, so it is invertible on $\mathfrak{su}(n)$. To make sense of $\Delta_{n}^{-1}w$” for $w\in\mathfrak{u}(n)$ more generally—as we will need to do at the internal stages—we can do one of two equivalent things:
  • Take the Moore–Penrose pseudoinverse, i.e., $\Delta_{n}^{-1}\bigl(w-\frac{1}{n}(\operatorname{tr}w)I\bigr)$.
  • View $\Delta_{n}$ as an invertible operator on the quotient $\mathfrak{pu}(n)=\mathfrak{u}(n)/i\mathbb{R}I$, as in Modin and Viviani [20, Section 3], and apply $\Delta_{n}^{-1}$ to the equivalence class of $w$. The representative of this equivalence class in $\mathfrak{su}(n)$ is exactly $w-\frac{1}{n}(\operatorname{tr}w)I$, as above.
Since the Lie bracket annihilates multiples of the identity, the term $\Delta_{n}^{-1}w,w]\in\mathfrak{su}(n)$ is thus well-defined for $w\in\mathfrak{u}(n)$, and we have $g\colon\mathfrak{u}(n)\rightarrow\mathfrak{su}(n)$ as needed.
It then follows from the preceding discussion that we can apply the method (4.4) to the system (4.5), where $L$ is skew-Hermitian, $P$ is Hermitian, and $M=N$, and this method is Lie–Poisson in the inviscid case $\nu=0$. Theorem 2.13 and Proposition 4.4 then give the following result on dissipation of energy $\frac{1}{2}\langle\Delta_{n}^{-1}w,w\rangle$ and enstrophy $\frac{1}{2}\lVert w\rVert^{2}$.
Corollary 4.6.
For the method (4.4) applied to the Navier–Stokes–Zeitlin system (4.5), energy and enstrophy are dissipated monotonically up to $\mathcal{O}(h^{3})$ when $b_{i}\geq 0$ for all $i$, and enstrophy is conserved exactly in the inviscid case $\nu=0$.
Proof.
Calculating $G^{\prime}(w)g(w)$ when $G$ is energy, we get
$-\bigl\langle\Delta^{-1}_{n}w,[\Delta_{n}^{-1}w,w]+\nu\Delta_{n}w\bigr\rangle$ $=-\nu\langle\Delta^{-1}_{n}w,\Delta_{n}w\rangle$
$=-\nu\langle\Delta_{n}\Delta_{n}^{-1}w,w\rangle$
$=-\nu\bigl\lVert w-\tfrac{1}{n}(\operatorname{tr}w)I\bigr\rVert^{2}$
$\leq 0,$
where the last equality uses the fact that the composition of $\Delta_{n}$ with its pseudoinverse is the orthogonal projection onto $\mathfrak{su}(n)$. For enstrophy, we get
$\bigl\langle w,[\Delta_{n}^{-1}w,w]+\nu\Delta_{n}w\bigr\rangle=\nu\langle w,\Delta_{n}w\rangle\leq 0,$
since $\Delta_{n}$ is negative-semidefinite. Hence, dissipation of both quantities up to $\mathcal{O}(h^{3})$ follows from Theorem 2.13. In the inviscid case $\nu=0$, we have $P=0$, so exact numerical conservation of enstrophy follows from Proposition 4.4. ∎
Acknowledgments
The authors wish to thank Klas Modin for helpful conversations and generous feedback during the early stages of this project. We also would like to thank the Isaac Newton Institute for Mathematical Sciences for support and hospitality during the program “Geometry, compatibility and structure preservation in computational differential equations,” supported by EPSRC grant number EP/R014604/1, as this paper grew out of conversations begun there. Ari Stern was supported by the National Science Foundation under Grant No. DMS-2208551.